### v2

# Degree-Weighted DeGroot Learning

Abstract
We examine belief formation in large stochastic networks under a degree-weighted DeGroot learning process. In many social settings, agents may place more trust in well-connected individuals or, conversely, discount their influence. Existing research on random networks primarily assumes that agents assign equal weights to their neighbors’ opinions. We relax this assumption by allowing agents to weight neighbors based on their degree. Our main technical contribution is to derive the asymptotic properties of learning outcomes, particularly the consensus belief and convergence speed, under this degree-weighted framework. Using this result, we analyze how the weighting rule affects consensus beliefs, societal wisdom, and convergence speed. We find that a more popularity-favoring rule—i.e., assigning greater weight to higher-degree neighbors—harms wisdom but has a non-monotonic effect on convergence speed. Whether it accelerates or slows convergence depends on the diversity of views within high- and low-degree groups. This highlights a potential tradeoff between faster convergence and an unwise consensus in networks where agents favor highly connected neighbors.
1 Introduction
The DeGroot model of social learning in networks has been instrumental in studying societal belief dynamics, such as whether society reaches consensus (DeGroot, 1974), agents’ network position and their long-run influence on consensus (DeMarzo, Vayanos and Zwiebel, 2003), the “wisdom of the crowd” phenomenon (Acemoglu, Ozdaglar and ParandehGheibi, 2010; Golub and Jackson, 2010), and the speed of belief convergence (Golub and Jackson, 2012). In a typical DeGroot learning model, agents repeatedly aggregate the opinions of their neighbors according to fixed weights over time. To gain tractability—especially when examining learning dynamics in large networks—agents are often assumed to assign equal weight to the opinions of all their neighbors (Golub and Jackson, 2012). However, the equal-weighting DeGroot model, while a useful benchmark, lacks certain empirically relevant elements. For example, it cannot capture attitudes toward popularity, where society may overweight the opinions of highly connected agents, believing that these individuals are exposed to a larger volume of information. Conversely, equal weighting fails to capture sentiment resistant to popularity, where society may underweight the opinions of highly connected agents, perceiving them as having more redundant or less trustworthy information sources. [^1] In practice, platforms may intentionally downweight highly popular users to provide “less-prominent” users with more visibility. For example, Facebook’s News Feed algorithm prioritizes content from friends, family, and close connections over posts from popular pages or celebrities. Similarly, Amazon’s initiatives, such as Amazon Launchpad and A+ Content, support new sellers and startups by offering enhanced product descriptions and media, thereby increasing their exposure and credibility. This approach improves the chances for smaller sellers to compete against established brands. [^2] The association between degree and low weighting does not need to be causal for the theory to hold. A high degree alone does not necessarily justify assigning lower weight to someone’s influence. Instead, it may be that individuals with certain characteristics (e.g., public figures with broad but less specialized appeal) tend to accumulate a high number of in-links, leading others to assign lower weight to their opinions. We thank the anonymous referee for suggesting this point.
This paper develops a degree-weighted DeGroot model in which agents assign weights to their neighbors’ opinions based on their degree. We allow the weight to increase with the degree of a neighbor, capturing the notion of popularity-favoring; conversely, a decreasing weight reflects the sentiment of popularity-resistance. [^3] When the weight is independent of degree, the model reduces to the classical equal-weighting case. Our main contribution is an asymptotic theorem that establishes that the learning outcomes (consensus belief and convergence speed) in our degree-weighted learning framework can be approximated using the “expected” learning matrix. [^4] Previous works, with Golub and Jackson (2012), Dasaratha (2020) and Mostagir and Siderius (2021) being the closest, do not address the convergence of highly non-linear terms arising from degree-dependent weighting. Building on this result, we extend the standard equal-weighting DeGroot learning framework to degree-weighted scenarios, introducing a new dimension of comparative statics and an additional degree of freedom in degree dependence, which was previously difficult to explore.
Key Takeaway. Our degree-weighted learning model reveals that degree dependence shapes both societal wisdom and the speed of belief convergence, highlighting a fundamental trade-off between the two. Stronger degree dependence amplifies the influence of high-degree individuals, potentially undermining wisdom and exhibiting a largely monotonic effect. Its impact on convergence speed, however, is non-monotonic: whether greater degree dependence accelerates convergence depends on the diversity of views within high-degree groups. These forces suggest a potential trade-off—greater degree dependence may lead to faster convergence, but at the cost of societal wisdom.
Wisdom. By fully characterizing agents’ influence—i.e., how their initial opinions shape the consensus belief—and the consensus belief itself, we apply the necessary and sufficient condition for a society to be wise, defined as the consensus belief uncovering the true state, as established in Golub and Jackson (2010). We find that degree-dependent weighting can shift a society between wisdom and its absence.
More specifically, we show that when agents’ initial beliefs equal the true state plus i.i.d. noise, degree dependence has a monotonic effect on societal wisdom. A high degree of popularity-favoring—meaning agents assign greater weight to high-degree neighbors—grants high-degree agents non-vanishing influence and renders the society unwise.
Convergence Speed. The degree dependence can have a highly non-monotonic impact on the speed at which societal beliefs reach consensus. We find that when agents exhibit a greater degree of popularity-favoring, society does not necessarily reach consensus faster. Whether increasing the degree of popularity-favoring accelerates or slows convergence depends on the diversity of views within high-degree and low-degree groups. When low-degree groups outnumber high-degree groups and different groups potentially hold different views, increasing the degree of popularity-favoring may accelerate convergence; the reverse occurs when high-degree groups outnumber low-degree groups.
Tradeoff Between Wisdom and Convergence Speed. These forces can create a tradeoff between faster convergence and a “wiser” society: while a more popularity-resistant society may foster greater wisdom, reaching it may take longer when many low-degree groups exist, each potentially holding different views. Similarly, a more popularity-favoring learning rule may accelerate belief convergence at the expense of wisdom.
Literature Review
Our paper relates to a substantial body of literature on DeGroot learning in social networks. DeGroot (1974) was among the first to examine naïve learning, where agents repeatedly aggregate their neighbors’ opinions according to a fixed set of weights, and identified conditions for belief convergence to consensus. [^5] Early contributions include French Jr (1956) and Harary (1959). For a recent survey on social learning and a discussion of the distinction between DeGroot and Bayesian models, see Golub and Sadler (2016). Subsequent work has extended the DeGroot paradigm to incorporate time-varying weights Chatterjee and Seneta (1977), initial biases or prejudices over agents’ prior beliefs Friedkin and Johnsen (1990), and opinion aggregation restricted to neighbors whose beliefs lie within a certain distance from one’s own Hegselmann and Krause (2002), among other variations. DeMarzo, Vayanos and Zwiebel (2003) later revisited the DeGroot model in an economic context, grounding naïve learning in a quasi-Bayesian framework and highlighting the role of persuasion bias in belief aggregation, along with other important insights. Gao (2021) examined centrality-weighted opinion dynamics but did not analyze the impact of such degree dependence on wisdom and convergence speed. Moreover, while these studies assumed deterministic networks, we analyze the DeGroot model in large, stochastic networks.
Acemoglu, Ozdaglar and ParandehGheibi (2010) and Golub and Jackson (2010) studied belief convergence to the true state in stochastic networks. Acemoglu, Ozdaglar and ParandehGheibi (2010) analyzed how “forceful agents” and network structure influence the gap between the consensus belief and the true state, showing that this gap shrinks as beliefs converge faster. Golub and Jackson (2010) identified conditions for societal wisdom under DeGroot learning and found that wisdom is independent of convergence speed. Our framework complements this literature by demonstrating that a third factor—degree dependence—affects both wisdom and convergence speed, creating a tradeoff between the two.
Most relatedly, Golub and Jackson (2012) examined how network features affect convergence speed in large networks, finding that homophily slows belief convergence. They derived the large-network property of the second-largest eigenvalue in modulus (SLEM), which is inversely related to convergence speed, for equal-weighting learning matrices. We extend their analysis to degree-dependent weighting matrices.
Dasaratha (2020) provided asymptotic characterizations of eigenvector and Katz-Bonacich centralities, while Mostagir and Siderius (2021) extended this analysis to other centrality measures, showing that, with high probability, these measures approximate their values in a well-defined “average” network. Unlike their focus on the centrality measures for the adjacency matrix, our work investigates the asymptotic properties of the consensus belief and the second-largest eigenvalue in modulus (SLEM) of the learning matrix, which includes nonlinear components arising from degree-dependent weighting.
Cerreia-Vioglio, Corrao and Lanzani (2024) extended the analysis to nonlinear robust belief aggregators (as opposed to linear DeGroot learning) and identified conditions for convergence, consensus, and the wisdom of the crowd. They showed that society attains wisdom when the SLEM falls below a certain threshold. Their discussion of SLEM relied on the large-sample property result from Golub and Jackson (2012) and thus depended on the equal-weighting assumption.
Last but not least, our study is related to a large body of mathematical literature that examines the convergence and mixing time of Markov chains. We discuss our relation to this literature alongside our technical discussion in the paper. Most interestingly, Sinatra et al. (2011) found that constructing maximal-entropy random walks requires step probabilities to be proportional to a power of the degree of the target node, which is similar to the degree-dependent weighting in our paper.
Summary. Our key technical contribution is deriving the large-sample properties of the consensus belief and SLEM for degree-weighted learning matrices, where the closest related work is Golub and Jackson (2012), which assumed equal weighting. This approach is not only of theoretical interest but also of economic relevance, allowing us to extend existing theories on consensus, the wisdom of the crowd, and convergence speed to settings with degree-dependent weighting. The analysis provides an additional degree of freedom for comparative statics on degree dependence and reveals a potential tradeoff between convergence speed and societal wisdom.
2 Model: Networks and Degree-Weighted Learning
[…]
DeGroot Learning Model. In the classical DeGroot (1974) learning model, an agent’s belief such as the probability of global warming or the safety of a COVID vaccine, is a simple linear aggregation of her neighbors’ beliefs in the last period. In other words, the belief vector $\mathbf{b}(t)$ at time $t\in\mathbb{N}=\{0,1,\ldots\}$ is given by:
$\mathbf{b}(t)=\mathbf{T}^{t}\mathbf{b}_{0}\,,$
in which $\mathbf{b}_{0}\in\mathbb{R}^{n}$ with $\|\mathbf{b}_{0}\|_{2}=1$ is an initial belief vector and $\mathbf{T}$, the learning matrix, is an $n\times n$ row-stochastic matrix, with each $T_{ij}$ representing the weight that agent $i$ puts on the belief of agent $j$ [^7] The constraint $\|\mathbf{b}_{0}\|_{2}=1$ is a standardization. Any initial belief $\mathbf{b}_{0}$ can be expressed as $\mathbf{b}_{0}=\|\mathbf{b}_{0}\|_{2}\cdot\mathbf{b}_{0}/\|\mathbf{b}_{0}\|_{2}$. Therefore, this effectively allows for any choice of initial belief.
Equal-Weighting Benchmark. Typically, agents are assumed to assign equal weight to their neighbors (e.g., Golub and Jackson, 2010, 2012). Under this assumption, the learning matrix $\mathbf{T}$ takes the following form:
$T_{ij}=\frac{A_{ij}}{d_{i}(\mathbf{A})}.$
Degree-Weighted Updating. While equal-weighting serves as a convenient and tractable benchmark, it falls short in capturing some real-life patterns. For instance, people may assign greater weight to celebrity opinions, or conversely, discount the views of more popular neighbors for various reasons discussed earlier. These observations suggest a learning heuristic in which the updating process depends on the popularity, or degree, of each source.
To capture these phenomena, we introduce a degree-weighted learning matrix, T, as follows:
$T_{ij}=\frac{A_{ij}d_{j}^{\alpha}}{\sum\limits_{j^{\prime}}A_{ij^{\prime}}d_{j^{\prime}}^{\alpha}}\,,$ (1)
where $d_{j}=d_{j}(\mathbf{A})$ is agent $j$’s degree in $\mathbf{A}$ and $\alpha\in\mathbb{R}$ captures how a neighbor’s weight depends on their degree. [^8] We note two points here. First, we allow for more general functional forms of degree dependence, $\phi(\alpha,d)\colon\mathbb{R}^{2}\rightarrow\mathbb{R}$, as long as they satisfy certain regularity conditions. We refer readers to Appendix B.1 for a more detailed discussion. Second, we acknowledge empirical evidence (e.g., Breza, Chandrasekhar and Tahbaz-Salehi, 2018) suggesting that agents may have limited knowledge of network structures and, in particular, may not know their neighbors’ degrees exactly. Our theoretical results can, to some extent, accommodate this lack of knowledge. As we show in the Appendix, our theory does not require exact knowledge of a neighbor’s degree. Instead, it suffices that each person has an impression of the number of friends that other individuals have. More precisely, what is required is a nonlinear transformation of the degree, i.e., $\phi(\alpha,d_{i})$. While the $i$-th individual may not directly observe $d_{j}$, we assume that they can observe $\phi(\alpha,d_{j})$ based on their impression of the $j$-th individual. And we have the following cases:
  • $\alpha=0$ (equal-weighting): the benchmark case studied in the literature (e.g., Golub and Jackson, 2012);
  • $\alpha>0$ (popularity-favoring): weights increase with a neighbor’s degree;
  • $\alpha<0$ (popularity-resistant): weights decrease with a neighbor’s degree.
Random Networks: Stochastic Block Model. Although DeGroot learning with deterministic networks has been extensively studied in the literature, its properties under random networks are less well understood. Random networks, however, offer distinct appeal and advantages over deterministic ones. [^9] For example, researchers often face data limitations and rely on statistical models of network formation, where links are probabilistically determined (see Dasaratha, 2020 for more examples). Moreover, in real-world settings, agents may encounter uncertainty regarding whom they will be matched with, which is better modeled using random networks (as in Acemoglu, Ozdaglar and ParandehGheibi, 2010). We use the stochastic block model to generate random networks in our paper (see Holland, Laskey and Leinhardt, 1983 for a classical discussion and Lee and Wilkinson, 2019 for a recent survey on its development).
Figure 1: A Network Generated by Stochastic Block Models
Notes: A network generated by a stochastic block model with $3$ groups, $30$ agents in each. The linking probability between agents $i$ and $j$ is given by $p_{ij}=0.3$ if $i$ and $j$ belong to the same group, and $0.04$ otherwise.
Suppose there are $n$ agents divided into $m$ groups. Let $\mathbf{n}=(n_{1},\ldots,n_{m})$, with $n_{k}$ being the number of agents in group $k\in\{1,\ldots,m\}$. Let P be an $m\times m$ symmetric matrix, which takes the following form:
$\textbf{P}=\begin{bmatrix}p_{11}&p_{12}&\ldots&p_{1m}\\ p_{21}&p_{22}&\ldots&p_{2m}\\ \vdots&\vdots&\ddots&\vdots\\ p_{m1}&p_{m2}&\ldots&p_{mm}\end{bmatrix}\,.$ (2)
P represents the meeting probabilities for each group, where $p_{ij}>0$ denotes the linking probability within the same group if $i=j$, or between different groups if $i\neq j$ [^10] We assume $p_{ij}>0$ in order for the expected adjacency matrix (definition follows) to be strongly connected and aperiodic to ensure convergence to consensus (see Proposition 1).
Let $\mathbf{A}(\mathbf{P},\mathbf{n})$ denote a random matrix generated by the corresponding stochastic block model. The entry $A_{ij}$ represents the element in the $i$-th row and $j$-th column of $\mathbf{A}$. Each $A_{ij}$ is an independent Bernoulli random variable, where $A_{ij}\sim\mathrm{Bernoulli}(P_{kl})$, for $i\neq j$, and agents $i$ and $j$ belong to groups $k$ and $l$, respectively.
Let $\mathbf{R}={\rm I}\kern-1.79993pt{\rm E}\mathbf{A}$ be the corresponding expected adjacency matrix. The expected degree of a node $i$ in some group $k_{i}$ is
$d_{i}(\mathbf{R})=\sum_{l=1,\ldots,m}p_{k_{i}l}n_{l}.$
Together with any $\alpha$, we define the expected learning matrix, $\mathbf{T}^{*}(\mathbf{R},\alpha)$, [^11] We note here that the role of matrices $\mathbf{R}$ and $\mathbf{T}^{*}$ is purely instrumental. They do not represent the actual network or the updating process but instead provide a tractable approximation that facilitates the analysis of the expectation and limiting behavior of the opinion dynamics. such that
$T^{*}_{ij}=\frac{R_{ij}d_{j}^{\alpha}(\mathbf{R})}{\sum\limits_{j^{\prime}}R_{ij^{\prime}}d_{j^{\prime}}^{\alpha}(\mathbf{R})}\,\forall i,j.$ (3)
Lastly, we impose the following assumptions to establish our asymptotic results:
Assumption 1 (Density).
Let $\tau_{n}=\min\limits_{i}\frac{d_{i}(\mathbf{R})}{n}$, we assume
$\lim_{n\rightarrow\infty}\frac{\tau_{n}}{\sqrt{\frac{\log n}{n}}}=\infty\,.$ (4)
Assumption 2 (Comparable Densities).
$\frac{\max_{1\leq i,j\leq m}\{p_{ij}\}}{\min_{1\leq i,j\leq m}\{p_{ij}\}}<\infty\,.$ (5)
Assumption 1 requires that the graph cannot be too sparse; hence, the minimal degree should be significantly larger than $\sqrt{n\log n}$. Consequently, this ensures that, with probability tending to 1, there exists a path connecting any pair of nodes. Assumption 2 requires that the nonzero connection probabilities in the network are comparable. While interaction probabilities can vary significantly, their ratios must remain bounded. We need both assumptions to guarantee beliefs converge (to consensus) in the long run as networks grow as well as that the realized networks are “close enough” to expected networks (our Theorem 1 in the next section). [^12] Our Assumption 2 aligns with Condition (4) in Definition 3 of Golub and Jackson (2012). Assumption 1 differs from the Condition (1) in Definition 3 of Golub and Jackson (2012), as it is specifically needed to prove the concentration of nonlinear terms, such as $d_{j}^{\alpha}$. We do not require Condition (2) of Definition 3 in Golub and Jackson (2012) as we work with non-zero linking probabilities $p_{ij}>0$.
3 General Results
In this section, we examine the impact of the degree-weighted updating rule on learning outcomes—particularly the consensus belief and convergence speed—in the following sequence. We first analyze learning outcomes in expected networks, which are deterministic, before presenting our asymptotic results. While the deterministic case builds on well-established findings in the literature, our primary contribution lies in characterizing the asymptotic properties of learning outcomes in random networks.
3.1 Results for Deterministic and Expected Networks
We begin with necessary definitions. For basic network terminology, we refer readers to standard textbooks such as Jackson et al. (2008).
Definition 1.
A network is called strongly connected if there exists a path connecting any two nodes of the network. A network is aperiodic if there is no integer $k>1$ dividing the length of every cycle of the network.
Definition 2 (Belief Convergence).
Beliefs are said to converge if, for any initial belief profile $b(0)$, the sequence $b(t)$ satisfies
$\lim_{t\to\infty}b_{i}(t)\text{ exists for all }i.$
Definition 3 (Consensus).
The network reaches consensus if beliefs converge to a common value, i.e.,
$\lim_{t\to\infty}b_{i}(t)=b^{*}\quad\forall i.$
where $b^{*}$ is a constant independent of $i$.
Definition 4 (Convergence Speed).
Denote $\mathbf{T}^{\infty}=\lim_{t\rightarrow\infty}\mathbf{T}^{t}$. Then, at time $t$, the convergence speed is measured by
$\max\limits_{\|\mathbf{b}_{0}\|_{2}=1}\|(\mathbf{T}^{t}-\mathbf{T}^{\infty})\mathbf{b}_{0}\|_{2},$ (6)
where $\|\cdot\|_{2}$ denotes the Euclidean norm. This expression quantifies the largest possible deviation between beliefs at time $t$ and the limiting belief, across all initial beliefs whose Euclidean norm is standardized to 1. A larger value indicates slower convergence.
Standard Results in the Literature.
The first well-established result in the literature is that if a network is strongly connected and aperiodic, long-run beliefs not only converge but also reach consensus (see DeGroot, 1974; Meyer, 2000; specifically, Proposition 6 in Golub and Sadler, 2017 for formal proofs and further discussion).
The second standard result in the literature is that the convergence speed of societal beliefs can usually be bound by the second largest eigenvalue in modulus (SLEM) of the learning matrix (see Montenegro, Tetali et al., 2006; Levin and Peres, 2017; see also Proposition 8 in Golub and Sadler, 2017 for a related discussion).
Consensus and Convergence Speed in Degree-Weighted Learning.
[…]
The adaptation of the convergence speed result to our setting is more intricate than in the previous case. Specifically, establishing the inverse relationship between the second largest eigenvalue in modulus (SLEM) and the convergence speed often relies on spectral decomposition, which assumes a diagonalizable learning matrix. A sufficient condition for a real matrix to be diagonalizable is symmetry. However, due to the degree-dependent weighting in our setting, the learning matrix is generally not symmetric. To address this, we show that it is similar to a symmetric matrix, [^14] See Lemma 1 in Appendix B.2.1 allowing us to apply the standard result linking the SLEM to the convergence speed.
We summarize the learning outcomes—particularly the consensus belief and convergence speed for any deterministic network and the expected learning matrix as a special case—in the following proposition.
Proposition 1 (Consensus and Convergence Speed for Deterministic and Expected Degree-Weighted Matrices).
Given any strongly connected and aperiodic adjacency matrix $\mathbf{A}$ with $A_{ij}>0$, [^15] The expected adjacency matrix $\mathbf{R}=\mathbb{E}[\mathbf{A}]$, generated by the stochastic block model, also satisfies this condition. Therefore, in this Proposition, we can replace the matrix $\mathbf{A}$ with $\mathbf{R}$ and $\mathbf{T}$ with $\mathbf{T}^{*}$ as special cases. the degree dependence $\alpha\in\mathbb{R}$, and the corresponding degree-weighted learning matrix $\mathbf{T}=\mathbf{T}(\mathbf{A};\alpha)$ as defined in Equation (3), we have:
  1. 1.
    Beliefs converge to a consensus $b(\infty)=(s_{1},\ldots,s_{n})\mathbf{b}_{0}$, where agent $i$’s influence $s_{i}$ is
    $s_{i}=\frac{\sum\limits_{j}A_{ij}\,d_{j}^{\alpha}(\mathbf{A})\,d_{i}^{\alpha}(\mathbf{A})}{\sum\limits_{i^{\prime},j}A_{i^{\prime}j}\,d_{j}^{\alpha}(\mathbf{A})\,d_{i^{\prime}}^{\alpha}(\mathbf{A})}.$ (7)
  2. 2.
    Let $|\lambda_{2}(\mathbf{T})|$ denote the second largest eigenvalue (in modulus) of $\mathbf{T}$ and let $\|\cdot\|_{2}$ denote the Euclidean norm. Then, at any time $t$,
    $\max\limits_{\|\mathbf{b}_{0}\|_{2}=1}\|(\mathbf{T}^{t}-\mathbf{T}^{\infty})\mathbf{b}_{0}\|_{2}=|\lambda_{2}(\mathbf{T})|^{t}.$ (8)
    Therefore, the convergence speed is inversely related to $|\lambda_{2}(\mathbf{T})|$.
Proof.
See Appendix B.3. ∎
We note a few things here. Our first result, which characterizes agent influence with degree-dependent weighting, helps us leverage previous results established in the literature that characterize conditions under which the consensus belief coincides with the true state (e.g., Lemma 1, p. 124 in Golub and Jackson, 2010). We devote Section 4.1 to a discussion of a wisdom-of-crowds type result.
Secondly, from the characterization of influence provided in Equation (7), one can see that when $\alpha=0$, we return to the equal-weighting case, and the agent’s influence is exactly as delivered in the classical literature (such as Theorem 6, p. 935 in DeMarzo, Vayanos and Zwiebel, 2003 and p. 122 in Golub and Jackson, 2010). Another special case is $\alpha=-1$. A brief calculation shows that, in this case, agents’ influence remains unequal; that is, their impact through degree is not fully offset by the popularity-resistant learning rule $\alpha=-1$ (see Section 4.1 for further discussion).
Lastly, by demonstrating that the second largest eigenvalue in modulus (SLEM) remains inversely related to the convergence speed in our case of degree-dependent weighting, we justify the examination of the asymptotic properties of the SLEM of a random learning matrix $\mathbf{T}$ and the comparative statics of $\alpha$ on the SLEM of the expected learning matrix $\mathbf{T^{*}}$.
3.2 Main Result: Random Networks and Asymptotic Results
In the previous section, we discussed standard results in the literature on consensus and the speed of belief convergence for deterministic networks. These results were extended to the expected adjacency and learning matrices in our case with degree-dependent weighting, $\mathbf{R}=\mathbb{E}[\mathbf{A}]$ and $\mathbf{T}^{*}=\mathbf{T}(\mathbf{R},\alpha)$, given any degree-dependence $\alpha$.
In this section, we focus on random networks and establish our key asymptotic results. In particular, we demonstrate that the learning outcomes corresponding to the realized networks, $\mathbf{A}$ and $\mathbf{T}$, are, in a probabilistic sense, sufficiently close to those corresponding to the expected networks, $\mathbf{R}$ and $\mathbf{T}^{*}$. [^16] In Appendix D, we show that our main results are also robust to further (minor) perturbations to the adjacency matrix $\mathbf{A}$.
Theorem 1.
Given any degree-dependence parameter $\alpha\in\mathbb{R}$ and a random network $\mathbf{A}(\mathbf{P},\mathbf{n})$, there exists a positive constant $\tilde{C}$, independent of $n$, such that for all $n>0$, the following results hold for the realized learning matrices $\mathbf{T}(\mathbf{A},\alpha)$ and the expected learning matrix $\mathbf{T}^{*}=\mathbf{T}(\mathbf{R},\alpha)$:
  1. 1.
    For influences,
    $\operatorname{\mathbb{P}}\Bigl{\{}\|\mathbf{s}(\mathbf{T})-\mathbf{s}(\mathbf{T}^{*})\|\geq\tilde{C}\sqrt{n\log n}\,(\min_{i}d_{i}(\mathbf{R}))^{-1}\Bigr{\}}\leq\frac{16}{n^{2}}\,.$
  2. 2.
    For learning speed,
    $\operatorname{\mathbb{P}}\Bigl{\{}\bigl{|}\lambda_{2}(\mathbf{T})-\lambda_{2}(\mathbf{T}^{*})\bigr{|}\geq\tilde{C}\frac{\sqrt{\log n}}{\tau_{n}\sqrt{n}}\Bigr{\}}\leq\frac{16}{n^{2}}\,.$
Sketch of Proof.
The formal proof is involved and is provided in Appendix B.4. Since Theorem 1 constitutes our primary technical contribution, we briefly outline the key ideas of the proof and indicate how our approach diverges from the existing literature.
The key steps in proving our asymptotic result are as follows:
  1. 1.
    To establish that $\mathbf{T}$ and $\mathbf{T}^{*}$ are similar to symmetric matrices, where
    $\mathbf{T}=\mathbf{D}_{1,\mathbf{A}}^{-1}\mathbf{A}\mathbf{D}_{2,\mathbf{A}},\quad\mathbf{T}^{*}=\mathbf{D}_{1,\mathbf{R}}^{-1}\mathbf{R}\mathbf{D}_{2,\mathbf{R}},$
    and $\mathbf{D}_{i,\mathbf{A}}$ and $\mathbf{D}_{i,\mathbf{R}}$ are diagonal matrices. [^17] See Lemma 1 in Appendix B.4, and Equations (15) and (18) in the Appendix for the above expressions.
  2. 2.
    To bound $|{\lambda_{2}}(\mathbf{T})-{\lambda_{2}}(\mathbf{T}^{*})|$, we apply Weyl’s inequality, which relies on the preceding result. This allows us to reduce the problem to establishing bounds for $\|\mathbf{A}-\mathbf{R}\|$, $\|\mathbf{D}_{1,\mathbf{A}}-\mathbf{D}_{1,\mathbf{R}}\|$, and $\|\mathbf{D}_{2,\mathbf{A}}-\mathbf{D}_{2,\mathbf{R}}\|$.
  3. 3.
    Bounding $\|\mathbf{A}-\mathbf{R}\|$ is standard in the literature (Golub and Jackson, 2012; Dasaratha, 2020; Mostagir and Siderius, 2021).
  4. 4.
    The difficulty lies in bounding $\|\mathbf{D}_{1,\mathbf{A}}-\mathbf{D}_{1,\mathbf{R}}\|$ and $\|\mathbf{D}_{2,\mathbf{A}}-\mathbf{D}_{2,\mathbf{R}}\|$, which are nonlinear in $\mathbf{A}$ and $\mathbf{R}$ in our degree-weighted framework.
  5. 5.
    Our approach: We use a leave-one-out technique (e.g., Abbe et al., 2020), replacing $d_{j}^{\alpha}(\mathbf{A})$ with $d_{j,-i}^{\alpha}(\mathbf{A})$ so that $A_{ij}$ is independent of $d_{j,-i}^{\alpha}(\mathbf{A})$, yielding $d_{j,-i}^{\alpha}(\mathbf{A})\approx d_{j}^{\alpha}(\mathbf{A})$ and thus
    $(\mathbf{D}_{1,\mathbf{A}})_{ii}\approx\sum_{j}A_{ij}\,d_{j,-i}^{\alpha}(\mathbf{A}).$
    Noting that, for fixed $i$, the collection $\{A_{ij}:j\neq i\}$ is independent of $\{d_{j,-i}(\mathbf{A}):j\neq i\}$, it follows that, conditional on $\{d_{j,-i}(\mathbf{A}):j\neq i\}$, the sum $\sum_{j\neq i}A_{ij}\,d_{j,-i}^{\alpha}(\mathbf{A})$ consists of independent random variables.
    This observation enables us to control the convergence rate of $(\mathbf{D}_{1,\mathbf{A}})_{ii}$ using standard concentration inequalities for sums of independent random variables. [^18] We refer the reader to Appendix B.4 for the complete proof.
Remark 1.
Proposition 1 and Theorem 1 remain valid when $\alpha$ depends on the individual $i$, i.e.,
$T_{ij}=\frac{A_{ij}d_{j}^{\alpha_{j}}}{\sum\limits_{j^{\prime}}A_{ij^{\prime}}d_{j^{\prime}}^{\alpha_{j^{\prime}}}},\quad T^{*}_{ij}=\frac{R_{ij}d_{j}^{\alpha_{j}}(\mathbf{R})}{\sum\limits_{j^{\prime}}R_{ij^{\prime}}d_{j^{\prime}}^{\alpha_{j^{\prime}}}(\mathbf{R})},\quad s_{i}=\frac{\sum\limits_{j}R_{ij}\,d_{j}^{\alpha_{j}}(\mathbf{R})d_{i}^{\alpha_{j}}(\mathbf{R})}{\sum\limits_{i^{\prime},j}R_{i^{\prime}j}d_{j}^{\alpha_{j}}(\mathbf{R})d_{i^{\prime}}^{\alpha_{j}}(\mathbf{R})},\quad\forall i,j,$
where $\max_{i,j}|\alpha_{i}-\alpha_{j}|\leq c$ for some positive constant $c$.
Proof.
See Appendix B.5. ∎
Theorem 1 lays the foundation for further discussion of economic insights in the degree‐weighted learning framework. In particular, to understand key learning outcomes, we are able to focus on the expected matrices rather than all possible realized networks. Together with Proposition 1 in the previous section, we know that influences (and later, the wisdom of a society) are captured by Equation (7), and that the convergence speed is inversely related to $|\lambda_{2}|$ of the expected learning matrix $\mathbf{T}^{*}$, as defined in Equation (8). In the next section, we provide a series of examples to highlight the impacts of degree‐weighted learning, as captured by $\alpha$, on societal wisdom, convergence speed, and a potential tradeoff between the two.
4 Applications
In this section, we apply the asymptotic results established above to three questions. First, we examine how the degree-dependent learning rule affects societal wisdom and find that its impact is largely monotonic. Next, we study its effect on the speed of learning, which displays a non-monotonic pattern. Lastly, we discuss a potential trade-off between societal wisdom and learning speed.
4.1 Influences and Wisdom of the Crowd
This section explores the concepts of influence and wisdom. [^19] We thank the editor and the anonymous referees for suggesting the discussion on wisdom. We begin by considering special cases of the degree-dependence parameter $\alpha$ to illustrate its impact on agents’ influences. Adopting the notion of wisdom introduced by Golub and Jackson (2010), we then show that variations in $\alpha$ are powerful enough to shift a society from wisdom to non-wisdom, or vice versa. Moreover, this effect is monotonic: increasing $\alpha$ toward a popularity-favoring learning rule leads to non-wisdom, while decreasing $\alpha$ toward a popularity-resistant rule promotes wisdom.
Influences With Special Cases of $\alpha$.
Recall that in Equation (7) of Proposition 1, we obtain the following characterization of the impact of $\alpha$ on influences for any deterministic network $\mathbf{A}$ that satisfies the imposed regularity conditions:
$s_{i}=\frac{\sum\limits_{j}A_{ij}\,d_{j}^{\alpha}d_{i}^{\alpha}}{\sum\limits_{i^{\prime},j}A_{i^{\prime}j}d_{j}^{\alpha}d_{i^{\prime}}^{\alpha}}.$
A notable special case is $\alpha=0$ (equal weighting). As shown in the literature (e.g., DeMarzo, Vayanos and Zwiebel, 2003 and Golub and Jackson, 2010), we have
$s_{i}=\frac{d_{i}}{\sum_{j}d_{j}},$
so that the influence of agent $i$ is proportional to his or her degree.
Our framework introduces additional interesting cases. For instance, the following example presents the case of $\alpha=-1$:
Example 1 (Influences when $\alpha=-1$).
When $\alpha=-1$, an agent $i$’s influence is given by
$s_{i}=\frac{\sum_{j}\frac{A_{ij}}{d_{i}d_{j}}}{\sum_{i^{\prime},j}\frac{A_{i^{\prime}j}}{d_{i^{\prime}}d_{j}}},$ (9)
which is inversely proportional to the (weighted) harmonic average of the neighbors’ degrees. When the neighbors’ degrees are independent, all agents share the same influence in expectation.
This specific degree dependence, $\alpha=-1$, approximately offsets the impact of degree heterogeneity, though it does not yield exactly equal influences.
Next, we highlight that the degree dependence, $\alpha$, via shaping the agents’ influences, can have more fundamental impacts on learning outcomes, such as determining whether the social learning process results in an unbiased aggregation of individual signals (i.e., the wisdom of the crowd).
Degree-Dependence and Wisdom.
We follow the definition of wisdom introduced in Golub and Jackson (2010). More specifically, suppose that agents’ initial beliefs are i.i.d. signals drawn from a distribution with mean $\theta$, which represents the true state of the world, and variance $\sigma^{2}>0$. To study networks in the large, we index the learning matrix by $n$, representing the size of the society. The sequence of learning matrices is then given by a sequence of $n$-by-$n$ interaction matrices. We denote the belief of agent $i$ in network $n$ at time $t$ by $b_{i}^{(t)}(n)$ and its limiting belief by $b_{i}^{(\infty)}(n)$. We say that a sequence $\bigl{(}T(n)\bigr{)}_{n=1}^{\infty}$ of learning matrices is wise if the limiting beliefs converge jointly in probability to the true state $\theta$.
Definition 5 (Wisdom).
The sequence $\bigl{(}T(n)\bigr{)}_{n=1}^{\infty}$ is wise if
$\lim_{n\to\infty}\max_{1\leq i\leq n}\Bigl{|}b_{i}^{(\infty)}(n)-\theta\Bigr{|}=0.$
Let $s(n)$ denote the influence vector corresponding to $T(n)$, as defined in Proposition 1. By applying Lemma 1 and Corollary 1 in Golub and Jackson (2010) (with our formulation for $s(n)$ replacing their $s(n)_{1}$), we obtain an explicit necessary and sufficient condition for a society to be wise when agents use a degree‐dependent updating rule.
Corollary 1 (Corollary 1 of Golub and Jackson (2010)).
A society is wise if and only if
$\lim_{n\to\infty}\max_{1\leq i\leq n}s_{i}(n)=0.$
Below, we present two examples that illustrate how a shift in the degree‐dependent updating parameter $\alpha$ can transform a society from unwise to wise, and vice versa; we then generalize the insights derived from these examples. All proofs are relegated to Appendix A.
Example 2 (Unwise to Wise).
Consider a sequence of stochastic block models with two types of agents: one high-degree agent and $n-1$ low-degree agents. The linking probabilities are given by $p_{11}=0$, $p_{12}=p_{21}=1$, and $p_{22}=\frac{1}{n-1}$, where $p_{ij}$ denotes the probability that an agent of type $i$ connects with an agent of type $j$. The resulting expected degrees are $d_{1}=n-1$ for the high-degree agent and $d_{2}=2$ for the low-degree agents. Let $s_{1}$ denote the influence of the high-degree agent.
Then, a shift from $\alpha=0$ to $-1$ changes the learning outcome from unwise to wise. As $n\to\infty$:
  • For $\alpha=0$, $s_{1}\to 1/3$, non-vanishing; the learning is unwise.
  • For $\alpha=-1$, $s_{1}\to 0$; the learning is wise. [^20] The influence of each low-degree agent is always negligible, as there are infinitely many of them.
Example 3 (Wise to Unwise).
Consider a revised version of the previous example with the following linking probabilities: $p_{11}=0$, $p_{12}=p_{21}=\frac{1}{\sqrt{n-1}}$, $p_{22}=\frac{1}{n-1}$. The resulting expected degrees are $d_{1}=\sqrt{n-1}$ for the high-degree agent and $d_{2}=1+\frac{1}{\sqrt{n-1}}$ for the low-degree agents.
Then, a shift from $\alpha=0$ to $2$ changes the learning outcome from wise to unwise. As $n\to\infty$:
  • For $\alpha=0$, $s_{1}\to 0$; the learning is wise.
  • For $\alpha=2$, $s_{1}\to\frac{1}{2}$; the learning is unwise.
The observation in Examples 2 and 3—that a sufficiently strong change in degree dependence $\alpha$ can shift a society from wise to unwise (and vice versa)—extends to more general settings. As long as the high-degree agents constitute a bounded group with unbounded expected degree, and the low-degree agents form an unbounded group with bounded expected degree, the effect of $\alpha$ on societal wisdom is monotonic. We refer to such networks as high-degree–low-degree networks and summarize the general impact in the following proposition. [^21] See Appendix A for a more detailed description of the general setting.
Proposition 2 (Monotonic Impact of $\alpha$ on Wisdom).
Given any high-degree–low-degree network as described above, there exists a threshold $\bar{\alpha}>-1$ such that the learning outcome is wise if and only if $\alpha<\bar{\alpha}$. As a special case, the learning outcome is always wise when $\alpha\leq-1$.
Proof.
See Appendix A. ∎
Thus, degree dependence $\alpha$ has a monotonic effect on societal wisdom: higher values of $\alpha$ reduce the likelihood that society is wise. This monotonicity arises because wisdom obtains if and only if every individual’s influence vanishes. Non-wisdom emerges when an increase in $\alpha$ grants a certain group non-vanishing influence. In our setting, a sufficiently large $\alpha$ may allow a bounded group—such as high-degree individuals (e.g., experts or media)—to retain influence, rendering the society unwise. The result reflects a structural asymmetry: high-degree agents form a bounded group with unbounded links, while low-degree agents form an unbounded group with bounded links, ensuring that only the former can sustain influence in large societies.
Key Takeaway on Societal Wisdom.
The above analysis shows that degree dependence $\alpha$ can significantly influence whether a society is wise or unwise. In networks with a bounded number of high-degree agents and infinitely many low-degree agents, the effect of $\alpha$ is monotonic: higher values of $\alpha$ grant high-degree agents non-vanishing influence and render the society unwise.
4.2 Degree-Dependence and Convergence Speed
In this section, we illustrate the effects of the degree-weighted learning heuristic, $\alpha$, on convergence speed. Our key finding is that in the high-degree–low-degree-group network that we examine, the impact of $\alpha$ on speed is non-monotonic. Whether a more popularity-favoring rule (larger $\alpha$) accelerates or slows down convergence depends on the diversity of views potentially held by each group.
More specifically, we study a model with two types of groups: high-degree and low-degree. There may be multiple groups of each type, with high-degree groups having a larger expected degree than low-degree groups. [^22] See Appendix C for a detailed description of the setting.
Convergence speed, measured by $|\lambda_{2}|$, can be fully characterized in the high-degree–low-degree model. [^23] See Proposition 6 in Appendix C. We show that the effect of $\alpha$ on $|\lambda_{2}|$ is highly non-monotonic. Moreover, the specific impact of $\alpha$ depends on the number of high- and low-degree groups, $m_{h}$ and $m_{l}$.
Example 4 (One High-Degree Group and One Low-Degree Group).
Let $m_{h}=m_{l}=1$. Then, there exists a threshold $\alpha^{*}$ such that convergence is slowest at $\alpha^{*}$. As $\alpha$ deviates from this threshold, $|\lambda_{2}|$ decreases, implying faster convergence. [^24] See Theorem 2 in Appendix C for the characterization of $\alpha^{*}$.
As illustrated in Example 4 (see the solid curve in Figure 2), changes in $\alpha$ lead to non-monotonic variations in learning speed. Convergence speed is not determined by a simple bias toward one group over the other; rather, the key determinant is the relative influence of the two groups. With a single high- and low-degree group, at most two dominant views emerge. When $\alpha$ is sufficiently large, the high-degree group dominates, leading to rapid convergence; likewise, as $\alpha\to-\infty$, the low-degree group dominates, also resulting in fast convergence. The slowest convergence occurs around an intermediate value $\alpha^{*}$, where the two groups exert relatively balanced influence.
Figure 2: Convergence Speed: One or Two High-Degree Groups and One Low-Degree Group
Notes: Convergence speed, measured by $|\lambda_{2}(\mathbf{T}^{*})|$, in degree dependence $\alpha$, for the following two cases:
One high-degree and one low-degree group (red solid): $n_{1}=10$ high-degree agents and $n_{2}=1000$ low-degree agents. The linking probabilities are $p_{11}=p_{22}=0.01$ within groups and $p_{12}=p_{21}=0.10$ across groups.
Two high-degree groups and one low-degree group (cyan dotted): $n_{1}=n_{2}=5$ high-degree agents and $n_{3}=1000$ low-degree agents. $p_{kk}=0.01$ within each group, $p_{k3}=p_{3k}=0.10$ for $k=1,2$ (between each high-degree group and the low-degree group), and $p_{12}=p_{21}=0.001$ between the two high-degree groups.
Example 5 (Two High-Degree Groups and One Low-Degree Group).
Let $m_{h}=2$ and $m_{l}=1$. The fastest convergence occurs as $\alpha\to-\infty$. For large positive values of $\alpha$, the pattern differs from Example 4, where there is only one high-degree group: convergence speed initially increases but then slows down as $\alpha\to+\infty$.
Examples 4 and 5 (see the dotted curve in Figure 2) exhibit markedly different patterns, despite having nearly identical degree distributions. [^25] The parameters are chosen so that in both Examples 4 and 5, there are 10 high-degree agents with $d_{h}\approx 100$ and 1000 low-degree agents with $d_{l}=11$. With multiple high-degree groups, diverse views may arise among them. Unlike the single high-degree group case in Example 4, the limit $\alpha\to+\infty$ does not lead to immediate convergence, as conflicting high-degree groups require time to reconcile their differing views. [^26] For general cases with multiple high- or low-degree groups and explicit characterizations of $|\lambda_{2}|$ and the threshold $\alpha^{*}$, see Proposition 6 and Theorem 2 in Appendix C.
Key Takeaway on Convergence Speed.
Our analysis reveals a non-monotonic impact of degree dependence on convergence speed. Greater popularity-favoring does not necessarily lead to faster consensus. Whether it accelerates or slows convergence depends on the diversity of views within high- and low-degree groups. When low-degree groups outnumber high-degree ones and different groups hold distinct views, greater popularity-favoring can accelerate convergence; the reverse occurs when high-degree groups outnumber low-degree groups.
4.3 Trade-off Between Convergence Speed and Wisdom
Our analysis of degree-weighted learning’s impact on societal wisdom and convergence speed reveals potential trade-offs between the two, offering new economic insights. For instance, while a society that resists popularity bias may achieve greater wisdom, reaching consensus can take longer when many low-degree groups exist, each potentially holding different views. Conversely, a more popularity-favoring learning rule may accelerate belief convergence at the cost of wisdom.
We summarize the effects of $\alpha$ on wisdom and convergence speed across different cases in the following table:
Table 1: Trade-off between Wisdom and Speed under Different Popularity Conditions
# Groups Degree Dependence
High Low Popularity-resistant ($\alpha\to-\infty$) Popularity-favoring ($\alpha\to\infty$)
1 1 fast, wise fast, unwise
multiple 1 slow, wise slow, unwise
1 multiple slow, wise fast, unwise
5 Discussion on the Symmetry Assumption
Before concluding, we discuss the symmetry assumption imposed on the adjacency matrix. Although both directed and undirected networks are natural in a learning setting, we focus on undirected networks because they are required for using the second largest eigenvalue in modulus (SLEM) to bound convergence speed and, more importantly, to derive our asymptotic result in this context.
Why Symmetry is Needed.
The derivation of the inverse relationship between SLEM and convergence speed requires employing spectral decomposition, which generally necessitates that the learning matrix be symmetric (or similar to a symmetric matrix). To prove the asymptotic result, we apply Weyl’s inequalities, which require the matrices to be Hermitian. Moreover, for real matrices, symmetry is both a necessary and sufficient condition for being Hermitian. These technical requirements are the key reasons why we focus on undirected networks.
What Happens When the Network is Asymmetric.
For non-symmetric matrices, there is no general eigenvalue inequality comparable to Weyl’s inequalities that can effectively bound the difference between $\mathbf{T}$ and $\mathbf{T}^{*}$. The convergence of the high-dimensional random matrix $\mathbf{T}^{\infty}$ (which is related to a non-symmetric matrix $\mathbf{A}$) is complex, as we do not know whether $\mathbf{T}$ is diagonalizable. We conjecture that our results still hold for diagonalizable $\mathbf{T}$.
Simulations.
Consider a stochastic block model with two groups, $n_{1}=500$ and $n_{2}=300$, and the linking matrix $\textbf{P}^{\prime}$ given by
$\mathbf{P}^{\prime}=\begin{pmatrix}0.3&0.2+a\\ 0.2-a&0.3\end{pmatrix},$ (10)
where $a\in[0,0.2]$ captures the level of asymmetry. When $a=0$, we recover the benchmark of symmetric matrices.
Figure 3a plots convergence speeds as a function of the degree dependence $\alpha$. For small asymmetry $a$, the convergence speed curves are close to the symmetric benchmark ($a=0$). When $a$ is large (e.g., $a=0.2$), the behavior differs significantly: $|\lambda_{2}|$ approaches 1 for high degree dependence $\alpha$, indicating a failure of convergence.
[…]
(a) Learning speed, $|\lambda_{2}|$.
(b) Convergence of $\mathbf{T}^{t}$, $2\log\|\mathbf{T}^{t}-(\mathbf{Q}^{-1})_{,1}\mathbf{Q}_{1,}\|_{F}$.
Figure 3: Learning Speed and Convergence of $\mathbf{T}^{t}$ With Various Levels of Asymmetry, $a$.
Notes: There are two groups with group sizes $n_{1}=500$ and $n_{2}=300$; linking probabilities are $p_{11}=p_{22}=0.3$, $p_{12}=0.2+a$, $p_{21}=0.2-a$, and $a$ captures the level of asymmetry.
Key Takeaways from the Simulations.
These simulations convey two key messages.
  • A slight departure from symmetry leads to similar learning outcomes, suggesting that the symmetric case is not a knife-edge condition and that our findings exhibit at least local robustness (see Figure 3 for relatively small $a$).
  • A large departure from symmetry, however, introduces substantial differences and challenges. For instance, in symmetric matrices, all eigenvalues are real, and $|\lambda_{2}|$ is strictly less than one, whereas in asymmetric matrices, eigenvalues may be complex, and $|\lambda_{2}|$ can approach one, indicating extremely slow convergence (see Figure 3 for $a=0.18$ or $0.2$).
6 Conclusion
Our views are shaped by neighbors in social networks. How we weigh their opinions affects the duration of disagreement and learning effectiveness in a society. To address these questions, we propose a degree-weighted DeGroot learning model, in which an agent’s weighting of neighbors’ opinions is determined by the neighbors’ popularity (degree). We examine how degree dependence affects consensus beliefs, societal wisdom, and the speed of belief convergence.
Our main technical contribution is establishing an asymptotic theory of learning outcomes (consensus beliefs and convergence speeds) for degree-weighted learning matrices in random networks. The central economic insight from this framework is that degree dependence shapes both convergence speed and societal wisdom, revealing a trade-off between the two: while a more popularity-favoring rule (higher degree dependence) may accelerate convergence when views among high-degree groups are relatively unified, it may hinder society’s ability to learn the true state.
These findings have broader implications for how societies process information with opinion leaders and centralized sources. Future research could extend the analysis to dynamic weighting rules, complex networks, or strategic belief updating. Understanding these mechanisms is crucial for designing information environments that balance rapid consensus with accurate collective learning.
Appendices
Appendix A Wisdom: General Settings and Related Proofs
A.1 General Setting for the Discussion on Wisdom
Consider a sequence of societies indexed by the population size $n$, with $n_{h}$ high-degree agents, $n_{l}=n-n_{h}$ low-degree agents, and a linking probability $q(n)$ for each pair consisting of one high-degree agent and one low-degree agent. We assume the following:
  1. 1.
    The group of high-degree agents remains bounded: $n_{h}=O(1)<\infty$, ensuring that the low-degree agents constitute an unbounded group with $n_{l}=n-n_{h}=O(n)$.
  2. 2.
    Each low-degree agent has a bounded expected degree, i.e., $d_{l}=O(1)<\infty$.
  3. 3.
    Each high-degree agent has an unbounded expected degree, i.e., $d_{h}\sim n\,q(n)\to\infty$, given that $n_{h}$ is bounded.
Notice that we do not require either the high-degree or low-degree agents to be homogeneous—there may be multiple subgroups within each category. Moreover, the linking structures within each group can be arbitrary, as long as the above assumptions hold.
Proposition 3 (Wisdom).
Given any $\alpha$, a sequence of societies as described above is wise if and only if
$\left(nq(n)\right)^{\max\{\alpha+1,2\alpha\}}=o(n).$
Proof.
We drop $(n)$ from the notations when there is no ambiguity. Let $H$ and $L$ be the sets of high- and low-degree agents, respectively. We have $|H|=n_{1}$ and $|L|=n_{2}=n-n_{1}$.
By Eq. (7) in Proposition 1, $i$’s influence $s_{i}=S_{i}/\sum_{i^{\prime}}S_{i^{\prime}}$, where $S_{i}\equiv\sum_{j}r_{ij}d_{i}^{\alpha}d_{j}^{\alpha}$, with $r_{ij}$ being the linking probability between $i$ and $j$. By assumption, $r_{hl}=q\text{ for }h\in H$ and $l\in L$.
It follows from the conditions that, $\forall h\in H$ and $l\in L$:
  • $n_{h}=O(1)$ (cond. 1) and $n_{l}=O(n)$;
  • $d_{l}-n_{h}\leq d_{l}=O(1)$ (cond. 2);
  • $d_{h}-n_{l}q=\sum_{h^{\prime}\in H}r_{ij}\leq n_{h}=O(1)$ and $d_{h}=O(n_{l}q)=O(nq)$; therefore
  • $d_{l}^{\alpha}=O(1)$ and $d_{h}=O\left((nq(n))^{\alpha}\right)$, $\forall\alpha$.
Now we are ready to characterize $S_{h}$ and $S_{l}$, as follows:
$S_{h}$ $=n_{l}qO\left((nq(n))^{\alpha}\right)+(d_{h}-n_{l}q)O\left((nq(n))^{2\alpha}\right)$
$=O\left((nq(n))^{\alpha+1}\right)+O\left((nq(n))^{2\alpha}\right)$
$=O\left((nq(n))^{\max\{\alpha+1,2\alpha\}}\right);$
$S_{l}$ $=n_{h}qO\left((nq(n))^{\alpha}\right)+(d_{l}-n_{h}q)O(1)$
$=O\left((nq(n))^{\alpha+1}\right)+O(n).$
High-degree agents’ influences:
$s_{h}=\frac{S_{h}}{\sum_{h\in H}S_{h}+\sum_{l\in L}S_{l}}=o(1)$$\iff$ $\frac{S_{h}}{nS_{l}}=o(1)$ $\iff(nq(n))^{\max\{\alpha+1,2\alpha\}}=o(n)$.
Low-degree agents’ influences:
$s_{l}=\frac{S_{l}}{n_{h}S_{h}+n_{l}S_{l}}=o(1)$ for sure.
Thus, the society is wise iff $(nq(n))^{\max\{\alpha+1,2\alpha\}}=o(n)$. ∎
A.2 Proofs Omitted in Section 4.1
Proofs of Example 2.
There is one high-degree agent with $d_{1}=n-1$ and $n-1$ low-degree agents with $d_{2}=2$.
When $\alpha=0$, the influence of the high-degree agent is given by:
$s_{1}=\frac{d_{1}}{d_{1}+(n-1)d_{2}}=\frac{n-1}{(n-1)+2(n-1)}=\frac{1}{3}.$
When $\alpha=-1$, the influence of the high-degree agent becomes:
$s_{1}=\frac{\frac{1}{d_{1}}}{\frac{1}{d_{1}}+(n-1)\left(\frac{1}{d_{2}}\left(\frac{1}{n-1}+\frac{1}{d_{2}}\right)\right)}\approx\frac{2}{n-1}\to 0,$
as $n\to\infty$.
[…]
Proof of Example 3.
In this model, there is one high-degree agent with $d_{1}=\sqrt{n-1}$ and $n-1$ low-degree agents with expected degree $d_{2}=1+\frac{1}{\sqrt{n-1}}$.
When $\alpha=0$, the influence of the high-degree agent is given by:
$s_{1}=\frac{d_{1}}{d_{1}+(n-1)d_{2}}=\frac{\sqrt{n-1}}{\sqrt{n-1}+(n-1)}\approx\frac{1}{\sqrt{n-1}}\to 0,$
as $n\to\infty$.
When $\alpha=2$, the influence of the high-degree agent is approximately:
$s_{1}\approx\frac{\sqrt{n-1}\,(\sqrt{n-1})^{2}}{\sqrt{n-1}\,(\sqrt{n-1})^{2}+(n-1)\Bigl{(}1+\sqrt{n-1}\Bigr{)}}.$
Since $(\sqrt{n-1})^{3}=(n-1)^{3/2}$ and $(n-1)(1+\sqrt{n-1})\sim(n-1)^{3/2}$ for large $n$, it follows that:
$s_{1}\to\frac{1}{2},\quad\text{as }n\to\infty.$
[…]
Proof of Proposition 2.
The result follows directly from Proposition 3. $\Box$
Appendix B Proofs for Results in Section 3
B.1 Preliminaries: Generalized Degree-Dependence Function
In this appendix, we prove that our results hold under a more general setting, where $d^{\alpha}$ is replaced by $\phi(\alpha,d)$. Here, $\phi(\alpha,d)$ satisfies Property 1–Property 3.
Property 1.
The function $\phi(\alpha,d)\in\mathcal{C}^{2}(\mathbb{R}\times[0,\infty))$ is nonnegative and satisfies $\phi(0,d)\equiv 1$ for all $d\in[0,\infty)$. Additionally, $\phi$ is monotonically increasing in $\alpha$ for $\alpha\in\mathbb{R}$, monotonically increasing in $d$ for $\alpha\in(0,\infty)$, and monotonically decreasing in $d$ for $\alpha\in(-\infty,0)$.
Property 2.
For any two degrees $d_{1}>d_{2}$, the ratio $\phi(\alpha,d_{2})/\phi(\alpha,d_{1})$ is strictly decreasing in $\alpha$:
$\frac{d}{d\alpha}\left(\frac{\phi(\alpha,d_{2})}{\phi(\alpha,d_{1})}\right)<0\,.$ (11)
Furthermore, we have the limits:
$\lim\limits_{\alpha\rightarrow\infty}\frac{\phi(\alpha,d_{2})}{\phi(\alpha,d_{1})}=\lim\limits_{\alpha\rightarrow-\infty}\frac{\phi(\alpha,d_{1})}{\phi(\alpha,d_{2})}=0.$ (12)
Property 3.
The function $\phi$ satisfies the following conditions:
$\limsup_{d\rightarrow\infty}\frac{\partial\phi(\alpha,d)/\partial d}{\phi(\alpha,d)/d}<\infty\,,$ (13)
and
$\limsup_{d\rightarrow\infty}\frac{\,\partial^{2}\phi(\alpha,d)/\partial d^{2}}{(\partial\phi(\alpha,d)/\partial d)/d}<\infty\,.$ (14)
Properties 1 and 2 establish fundamental requirements for $\phi$ to ensure the validity of our discussion. Property 3 consists of two technical conditions necessary for the concentration result derived later for random networks. Intuitively, Property 3 ensures that a small change in the degree $d$ does not lead to an excessive increase in $\phi$.
B.2 Proof Preliminaries
Throughout the appendix, various equalities and inequalities involve constants. We use the capital letter $C$ to denote such constants, which may vary across expressions but remain positive and independent of $n$.
B.2.1 $\mathbf{T}$ Similar to a Symmetric Matrix
As a preliminary step for the following proofs, we first establish that $\mathbf{T}$ is similar to a symmetric matrix.
Lemma 1.
The learning matrix $\mathbf{T}$, as defined in Equation (3), is similar to a symmetric matrix.
Proof.
The matrix $\mathbf{T}$ is given by
$\mathbf{T}=\mathbf{D}_{1,\mathbf{A}}^{-1}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}\,,$ (15)
where $\mathbf{D}_{1,\mathbf{A}}$ and $\mathbf{D}_{2,\mathbf{A}}$ are diagonal matrices with diagonal entries:
$(\mathbf{D}_{1,\mathbf{A}})_{ii}=\sum_{j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))\,,$
$(\mathbf{D}_{2,\mathbf{A}})_{ii}=\phi(\alpha,d_{i}(\mathbf{A}))\,.$
Define the diagonal matrix $\mathbf{S}$ as
$\mathbf{S}=\left(\mathbf{D}_{1,\mathbf{A}}\mathbf{D}_{2,\mathbf{A}}\right)^{-1/2}.$ (16)
Since the diagonal matrices $\mathbf{D}_{1,\mathbf{A}}$ and $\mathbf{D}_{2,\mathbf{A}}$ commute, we can rewrite $\mathbf{T}$ in terms of $\mathbf{S}$ as
$\mathbf{S}^{-1}\mathbf{T}\mathbf{S}=\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}.$ (17)
Since $\mathbf{A}$ is symmetric, the right-hand side is also symmetric, proving that $\mathbf{T}$ is similar to a symmetric matrix. ∎
A direct corollary of the previous result is that $\mathbf{T}^{*}$ can be expressed as
$\mathbf{T}^{*}=\mathbf{D}^{-1}_{1,\mathbf{R}}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}\,,$ (18)
where $\mathbf{R}=\mathbb{E}[\mathbf{A}]$, and $\mathbf{D}_{1,\mathbf{R}}$ and $\mathbf{D}_{2,\mathbf{R}}$ are diagonal matrices defined as
$(\mathbf{D}_{1,\mathbf{R}})_{ii}=\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))\,,$
$(\mathbf{D}_{2,\mathbf{R}})_{ii}=\phi(\alpha,d_{i}(\mathbf{R}))\,.$
B.3 Proof of Proposition 1
B.3.1 Proof of the First Part of Proposition 1
The following lemma, commonly known as the Perron-Frobenius Theorem, establishes that long-run beliefs converge to a consensus.
Lemma 2.
For a positive, stochastic matrix $\mathbf{T}$, the spectral radius $\rho(\mathbf{T})=1$ is a simple eigenvalue. Moreover, the limit $\mathbf{T}^{\infty}$ exists and is given by
$\mathbf{T}^{\infty}=\lambda_{1}\mathbf{U}_{1}=\mathbf{v}\mathbf{w}^{\top}\,,$
where $\mathbf{v}$ and $\mathbf{w}$ are the left and right eigenvectors of $\mathbf{T}$ corresponding to $\lambda_{1}$, normalized such that $\mathbf{w}^{\top}\mathbf{v}=1$.
The proof can be found in Meyer (2000).
Next, we introduce another lemma that will also be used in the proof.
Lemma 3.
Let $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ be similar $n\times n$ matrices such that
$\mathbf{M}_{2}=\mathbf{P}\mathbf{M}_{1}\mathbf{P}^{-1},$
where $\mathbf{P}$ is an invertible $n\times n$ matrix. If $\mathbf{v}$ is an eigenvector of $\mathbf{M}_{1}$ corresponding to the eigenvalue $\lambda$, then $\mathbf{P}\mathbf{v}$ is an eigenvector of $\mathbf{M}_{2}$ corresponding to $\lambda$.
Proof.
Since $\mathbf{v}$ is an eigenvector of $\mathbf{M}_{1}$ with eigenvalue $\lambda$, we have
$\mathbf{M}_{1}\mathbf{v}=\lambda\mathbf{v}.$
Multiplying both sides by $\mathbf{P}$ gives
$\mathbf{P}\mathbf{M}_{1}\mathbf{v}=\lambda\mathbf{P}\mathbf{v}.$
Using the similarity relation $\mathbf{M}_{2}=\mathbf{P}\mathbf{M}_{1}\mathbf{P}^{-1}$, we obtain
$\mathbf{M}_{2}(\mathbf{P}\mathbf{v})=\mathbf{P}\mathbf{M}_{1}\mathbf{P}^{-1}(\mathbf{P}\mathbf{v})=\mathbf{P}\mathbf{M}_{1}\mathbf{v}=\lambda\mathbf{P}\mathbf{v}.$
Thus, $\mathbf{P}\mathbf{v}$ is an eigenvector of $\mathbf{M}_{2}$ corresponding to $\lambda$, as desired. ∎
We seek to evaluate the limit:
$\mathbf{T}^{\infty}=\lim_{t\rightarrow\infty}\mathbf{T}^{t}=\lim_{t\rightarrow\infty}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{D}_{2,\mathbf{A}}^{-1/2}\left(\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\right)^{t}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{D}_{2,\mathbf{A}}^{1/2}.$ (19)
By Lemma 2, we know that the limit converges:
$\lim_{t\rightarrow\infty}\left(\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\right)^{t}=\lambda_{1}\mathbf{v}_{1}\mathbf{v}_{1}^{\top},$ (20)
where $\lambda_{1}=1$ is the largest eigenvalue in magnitude, and $\mathbf{v}_{1}$ is the corresponding unit eigenvector of the matrix $\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}$. Since the matrix in (20) is symmetric, its left and right eigenvectors are identical.
To determine $\mathbf{v}_{1}$, we apply Lemma 3. It is straightforward to verify that $\mathbf{D}_{1,\mathbf{A}}^{-1}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}$ has the eigenvector $\mathbf{e}=(1,\dots,1)^{\top}$ corresponding to eigenvalue $1$. Consequently, the matrix $\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}$ has the eigenvector
$\mathbf{w}_{1}=\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{e}.$
Normalizing $\mathbf{w}_{1}$ to obtain a unit vector, we divide by its Euclidean norm:
$\|\mathbf{w}_{1}\|_{2}=\left\|\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{e}\right\|_{2}=\left(\sum_{i,j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))\phi(\alpha,d_{i}(\mathbf{A}))\right)^{1/2}.$
Thus, the unit eigenvector is given by $\mathbf{v}_{1}=\mathbf{w}_{1}/\|\mathbf{w}_{1}\|_{2}$, and substituting into (19) yields
$\mathbf{T}^{\infty}$ $=\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{D}_{2,\mathbf{A}}^{-1/2}\left(\|\mathbf{w}_{1}\|_{2}^{-1}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{e}\right)\left(\|\mathbf{w}_{1}\|_{2}^{-1}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{e}\right)^{\top}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{D}_{2,\mathbf{A}}^{1/2}$
$=\|\mathbf{v}_{1}\|_{2}^{-2}\mathbf{E}_{n}\mathbf{D}_{1,\mathbf{A}}\mathbf{D}_{2,\mathbf{A}},$
where $\mathbf{E}_{n}$ is an $n\times n$ matrix with all entries equal to $1$. This gives the final expression:
$\mathbf{T}^{\infty}_{ij}=\frac{\sum_{j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))\phi(\alpha,d_{i}(\mathbf{A}))}{\sum_{i,j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))\phi(\alpha,d_{i}(\mathbf{A}))},\quad\forall i,j=1,\dots,n.$ (21)
$\blacksquare$
B.3.2 Proof of the Second Part of Proposition 1
The second part generalizes the result in Golub and Jackson (2012). For the proof, we establish an upper and a lower bound and show that they coincide. The main structure of the proof follows Golub and Jackson (2012), with modifications to accommodate our matrix $\mathbf{T}$ and the norm $\|\cdot\|_{2}$.
A key step in this adaptation is applying the spectral theorem to decompose $\mathbf{T}$. Lemma 1 ensures that we can apply the spectral theorem for diagonalizable matrices Meyer (2000) to $\mathbf{T}$. The spectral decomposition of $\mathbf{T}$ is given by
$\mathbf{T}=\sum_{i=1}^{n}\lambda_{i}\mathbf{U}_{i}\,,$ (22)
where $\lambda_{i}$ are the eigenvalues of $\mathbf{T}$, ordered in decreasing magnitude, and $\mathbf{U}_{i}$ are the orthogonal projections onto the corresponding eigenspaces of $\mathbf{T}$.
Upper Bound
We aim to establish an upper bound for the distance $\left\|\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}\right\|_{2}$ by applying the spectral decomposition of $\mathbf{T}$ from (22). Note that
$\mathbf{T}^{t}-\mathbf{T}^{\infty}=\sum_{i=2}^{n}\lambda_{i}^{t}\mathbf{U}_{i}\,.$ (23)
The summation starts from $i=2$ due to Lemma 2, which states that the largest eigenvalue of $\mathbf{T}$ is $1$, while all other eigenvalues have magnitudes strictly less than $1$. The term corresponding to $i=1$ cancels with $\mathbf{T}^{\infty}$ since
$\mathbf{U}_{1}=(\mathbf{v}\mathbf{w}^{\top})^{t}=\mathbf{v}(\mathbf{w}^{\top}\mathbf{v})^{t-1}\mathbf{w}^{\top}=\mathbf{v}\mathbf{w}^{\top}=\mathbf{T}^{\infty},\quad\text{for }t\in\mathbb{N}.$
Applying (23), we obtain
$\left\|\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}\right\|_{2}^{2}$ $=\left\lVert\sum_{i=2}^{n}\lambda_{i}^{t}\mathbf{U}_{i}\mathbf{b}\right\rVert^{2}_{2}$ (24)
$=\sum_{i=2}^{n}|\lambda_{i}|^{2t}\left\lVert\mathbf{U}_{i}\mathbf{b}\right\rVert^{2}_{2},\quad\quad(\mathbf{U}_{i}\mathbf{U}_{j}=0\text{ for }i\neq j)$
$\leq|\lambda_{2}|^{2t}\sum_{i=2}^{n}\left\lVert\mathbf{U}_{i}\mathbf{b}\right\rVert^{2}_{2}$
$=|\lambda_{2}|^{2t}\left\lVert\sum_{i=2}^{n}\mathbf{U}_{i}\mathbf{b}\right\rVert^{2}_{2}\,.\quad\quad(\mathbf{U}_{i}\mathbf{U}_{j}=0\text{ for }i\neq j)$
Define $\mathbf{U}=\sum_{i=2}^{n}\mathbf{U}_{i}$, which is a single orthogonal projection satisfying $\mathbf{U}=\mathbf{U}^{\top}=\mathbf{U}^{2}$. Then, for any vector $\mathbf{x}$,
$\|\mathbf{U}\mathbf{x}\|^{2}_{2}=\langle\mathbf{U}\mathbf{x},\mathbf{U}\mathbf{x}\rangle=\langle\mathbf{U}^{\top}\mathbf{U}\mathbf{x},\mathbf{x}\rangle=\langle\mathbf{U}\mathbf{x},\mathbf{x}\rangle\leq\|\mathbf{U}\mathbf{x}\|_{2}\|\mathbf{x}\|_{2}.$
The last inequality follows from the Cauchy-Schwarz inequality. Applying this to (24), we obtain
$\left\|\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}\right\|_{2}^{2}\leq|\lambda_{2}|^{2t}\|\mathbf{b}\|_{2}^{2}=|\lambda_{2}|^{2t}.$
Taking the square root on both sides yields the desired upper bound. $\blacksquare$
Lower Bound
To derive a lower bound, we consider a specific choice of $\mathbf{b}$. Let $\mathbf{b}$ be the eigenvector of $\mathbf{T}$ corresponding to its second largest eigenvalue $\lambda_{2}$ in magnitude, normalized such that $\|\mathbf{b}\|_{2}^{2}=1$. Then,
$\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}=\lambda_{2}^{\,t}\mathbf{b}.$
This follows from the fact that $\mathbf{U}_{i}\mathbf{b}=0$ for all $i\neq 2$. Consequently,
$\left\|\left(\mathbf{T}^{t}-\mathbf{T}^{\infty}\right)\mathbf{b}\right\|_{2}^{2}$ $=\left\lVert\lambda_{2}^{\,t}\mathbf{b}\right\rVert^{2}_{2}$
$=|\lambda_{2}|^{2t}.$
Taking the square root on both sides confirms that the lower bound matches the upper bound.
This completes the proof of Proposition 1. $\Box$
B.4 Proof of Theorem 1
B.4.1 Proof of the First Part of Theorem 1
Similar to Equation (19), we have
$(\mathbf{T}^{*})^{\infty}=\lim_{t\rightarrow\infty}(\mathbf{T}^{*})^{t}=\lim_{t\rightarrow\infty}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{D}_{2,\mathbf{R}}^{-1/2}\left(\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{R}_{1,\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\right)^{t}\mathbf{D}_{1,\mathbf{R}}^{1/2}\mathbf{D}_{2,\mathbf{R}}^{1/2}\,.$
Combining this with (19) and (20), we have
$\|\mathbf{s}(\mathbf{T})-\mathbf{s}(\mathbf{T}^{*})\|\leq\|\mathbf{T}^{\infty}-(\mathbf{T}^{*})^{\infty}\|=$ $\|\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{D}_{2,\mathbf{A}}^{-1/2}\mathbf{v}_{1}\mathbf{v}_{1}^{\top}\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{D}_{2,\mathbf{A}}^{1/2}-\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{D}_{2,\mathbf{R}}^{-1/2}\mathbf{v}^{*}_{1}\mathbf{v}_{1}^{*\top}\mathbf{D}_{1,\mathbf{R}}^{1/2}\mathbf{D}_{2,\mathbf{R}}^{1/2}\|$
$\leq$ $\|\mathbf{v}_{1}\mathbf{v}_{1}^{\top}-\mathbf{v}^{*}_{1}\mathbf{v}_{1}^{*\top}\|+\|\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{D}_{2,\mathbf{A}}^{-1/2}\mathbf{D}_{1,\mathbf{R}}^{1/2}\mathbf{D}_{2,\mathbf{R}}^{1/2}-\mathbf{I}\|$
$+\|\mathbf{D}_{1,\mathbf{A}}^{1/2}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{D}_{2,\mathbf{R}}^{-1/2}-\mathbf{I}\|,$ (25)
where $\mathbf{v}^{*}_{1}$ is the first eigenvector of $\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{R}_{1,\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}$.
In order to control the right hand side of the inequality, we introduce the following propositions and lemmas, with their proof presented at the end of the proof of Theorem 1.
Proposition 4.
$\|\mathbf{A}-\mathbf{R}\|\leq 3\sqrt{n\log n}\,,$
with probability at least $1-\frac{2}{n^{2}}$.
Proposition 5.
Under Assumptions (1)-(2),
$\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|\leq C\sqrt{n\log n}\left(\min_{i}d_{i}(\mathbf{R})\right)^{-3/2}\,,$ (26)
with probability at least $1-\frac{14}{n^{2}}$.
Lemma 4.
Let $K>1$ be a constant independent of $n$. Then, for $\alpha\geq 0$,
$\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))\leq C_{1}\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\,,$ (27)
and
$C_{2}\frac{\partial\phi}{\partial d}(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\frac{\partial\phi}{\partial d}(\alpha,K\min_{i}d_{i}(\mathbf{R}))\leq C_{3}\frac{\partial\phi}{\partial d}(\alpha,\min_{i}d_{i}(\mathbf{R}))\,,$ (28)
where $C_{1},C_{2},C_{3}$ are positive constants independent of $n$.
On the other hand, for $\alpha<0$,
$C_{1}\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))\leq\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\,,$ (29)
and
$C_{3}\frac{\partial\phi}{\partial d}(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\frac{\partial\phi}{\partial d}(\alpha,K\min_{i}d_{i}(\mathbf{R}))\leq C_{2}\frac{\partial\phi}{\partial d}(\alpha,\min_{i}d_{i}(\mathbf{R}))\,,$ (30)
Lemma 5.
For a fixed $j$,
$|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})|\leq\sqrt{2n\log n}\,,$ (31)
with probability at least $1-\frac{2}{n^{4}}$.
Moreover,
$\left|\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j}(\mathbf{R}))\right|\leq\sqrt{2n\log n}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,1})\right|\,,$ (32)
for some $c_{j,1}\in\left[d_{j}(\mathbf{R})-\sqrt{2n\log n},d_{j}(\mathbf{R})+\sqrt{2n\log n}\right]$ for each $j=1,\ldots,n$, with probability at least $1-\frac{2}{n^{4}}$.
We first control the second term on the right-hand side of (B.4.1). By Lemmas 45 and Proposition 5, with probability at least $1-\frac{15}{n^{2}}$, we obtain
$\|\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{D}_{2,\mathbf{A}}^{-1/2}\mathbf{D}_{1,\mathbf{R}}^{1/2}\mathbf{D}_{2,\mathbf{R}}^{1/2}-\mathbf{I}\|$ $\leq\|\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{D}_{\mathbf{R}}^{1/2}-\mathbf{I}\|\|\mathbf{D}_{2,\mathbf{R}}\mathbf{D}_{2,\mathbf{A}}^{-1}\|+\|\mathbf{D}_{2,\mathbf{R}}\mathbf{D}_{2,\mathbf{A}}^{-1}-\mathbf{I}\|$ (33)
$\leq C\sqrt{n\log n}(\min_{i}d_{i}(\mathbf{R}))^{-1},$
where $C$ is a positive constant. A similar inequality holds for the third term on the right-hand side of (B.4.1).
For the first term on the right-hand side of (B.4.1), using a similar argument as in the proof of (33) and applying the Davis-Kahan theorem along with Propositions 45 and Lemmas 45, with probability at least $1-\frac{15}{n^{2}}$, we obtain
$\|\mathbf{v}_{1}\mathbf{v}_{1}^{\top}-\mathbf{v}^{*}_{1}\mathbf{v}_{1}^{*\top}\|\leq C\|\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}-\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\|,$ (34)
which further satisfies
$\|\mathbf{v}_{1}\mathbf{v}_{1}^{\top}-\mathbf{v}^{*}_{1}\mathbf{v}_{1}^{*\top}\|\leq C\sqrt{n\log n}(\min_{i}d_{i}(\mathbf{R}))^{-1},$ (35)
where $C$ is a positive constant.
This completes the proof of the first part of Theorem 1. $\blacksquare$
B.4.2 Proof of the Second Part of Theorem 1
Recall that $\mathbf{R}={\rm I}\kern-1.79993pt{\rm E}\mathbf{A}$. From the definitions of the matrices in Equations (15) and (18), and by Lemma 3, comparing the eigenvalues $\lambda_{2}(\mathbf{T})$ and $\lambda_{2}(\mathbf{T}^{*})$ reduces to comparing the eigenvalues of
$\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{2,\mathbf{A}}^{1/2}\mathbf{D}_{1,\mathbf{A}}^{-1/2}$
and
$\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}^{1/2}\mathbf{D}_{1,\mathbf{R}}^{-1/2},$
since these matrices are similar to $\mathbf{T}$ and $\mathbf{T}^{*}$, respectively.
For simplicity, define
$\mathbf{D}_{\mathbf{A}}=\mathbf{D}_{2,\mathbf{A}}^{-1}\mathbf{D}_{1,\mathbf{A}}\quad\text{and}\quad\mathbf{D}_{\mathbf{R}}=\mathbf{D}_{2,\mathbf{R}}^{-1}\mathbf{D}_{1,\mathbf{R}}.$
By Weyl’s inequality, we obtain
$|\lambda_{2}(\mathbf{T})-\lambda_{2}(\mathbf{T}^{*})|\leq\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{\mathbf{R}}^{-1/2}\right\rVert,$ (36)
[…]
Under Assumptions (1) and (2), the expected degrees satisfy:
$1<\liminf_{n}\frac{\min_{i}d_{i}(\mathbf{R})}{\max_{i}d_{i}(\mathbf{R})}\leq\limsup_{n}\frac{\max_{i}d_{i}(\mathbf{R})}{\min_{i}d_{i}(\mathbf{R})}<\infty,$
which implies the existence of a positive constant $C$ such that for all $n$,
$\min_{i}d_{i}(\mathbf{R})<\max_{i}d_{i}(\mathbf{R})<C\min_{i}d_{i}(\mathbf{R}).$ (37)
Furthermore, we have
$\min_{i}(\mathbf{D}_{\mathbf{R}})_{ii}=\min_{i}\frac{\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))}{\phi(\alpha,d_{i}(\mathbf{R}))}=\min_{i}\sum_{j}R_{ij}\frac{\phi(\alpha,d_{j}(\mathbf{R}))}{\phi(\alpha,d_{i}(\mathbf{R}))}\geq\min_{i}d_{i}(\mathbf{R}).$ (38)
First, note that the spectral norm in (36) can be decomposed as:
$\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{\mathbf{R}}^{-1/2}\right\rVert$ (39)
$\leq$ $\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{A}}^{-1/2}\right\rVert+\left\lVert\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{R}}^{-1/2}\right\rVert$
$+\left\lVert\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{R}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{\mathbf{R}}^{-1/2}\right\rVert$
$\leq$ $\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|\cdot\|\mathbf{A}\|\cdot\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}\right\|+\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|\cdot\|\mathbf{A}\|\cdot\left\|\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|$
$+\|\mathbf{A}-\mathbf{R}\|\cdot\left\|\mathbf{D}_{\mathbf{R}}^{-1}\right\|.$
By the triangle inequality,
$\|\mathbf{A}\|\leq\|\mathbf{A}-\mathbf{R}\|+\|\mathbf{R}\|\,.$ (40)
The spectral norm is bounded above by the Frobenius norm:
$\|\mathbf{R}\|\leq\|\mathbf{R}\|_{F}=\sqrt{\sum_{i,j}R^{2}_{ij}}\leq Cn\max_{i,j}P_{ij}\leq C\min_{j}d_{j}(\mathbf{R})\,,$ (41)
for all $n>0$, where $\max_{i,j}P_{ij}$ is the largest entry of the matrix $\mathbf{P}$ in the stochastic block model.
Applying Proposition 4 along with Assumption 1, we obtain
$\|\mathbf{A}\|\leq C\min_{j}d_{j}(\mathbf{R})\,,$ (42)
with probability at least $1-\frac{2}{n^{2}}$. Similarly, by Proposition 5,
$\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\right\rVert\leq C\left(\min_{j}d_{j}(\mathbf{R})\right)^{-1/2}\,,$ (43)
with probability at least $1-\frac{14}{n^{2}}$, for sufficiently large $n$.
Conditioning on the event
$\left\{\|\mathbf{A}-\mathbf{R}\|\leq 3\sqrt{n\log n}\right\}\bigcap\left\{\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|\leq C\sqrt{n\log n}\left(\min_{i}d_{i}(\mathbf{R})\right)^{-3/2}\right\},$ (44)
which holds with probability at least $1-\frac{16}{n^{2}}$, we derive
$\left\lVert\mathbf{D}_{\mathbf{A}}^{-1/2}\mathbf{A}\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\mathbf{R}\mathbf{D}_{\mathbf{R}}^{-1/2}\right\rVert\leq C\frac{\sqrt{n\log n}}{\min_{i}d_{i}(\mathbf{R})}=C\frac{\sqrt{\log n}}{\tau_{n}\sqrt{n}}\,.$ (45)
This completes the proof for symmetric $\mathbf{A}$ and concludes our proof of Theorem 1. $\Box$
B.5 Supplementary Proofs of Theorem 1
Proof of Proposition 4
To bound $\|\mathbf{A}-\mathbf{R}\|$, we rewrite it as:
$\mathbf{A}-\mathbf{R}=\mathbf{Y}+\sum_{1\leq i<j\leq n}\mathbf{X}^{i,j},$
where
$\mathbf{X}^{i,j}=(\mathbf{A}-\mathbf{R})(\mathbf{e}_{i}\mathbf{e}_{j}^{\top}+\mathbf{e}_{j}\mathbf{e}_{i}^{\top}),$
$\mathbf{Y}=\operatorname{diag}(A_{11}-R_{11},\ldots,A_{nn}-R_{nn}),$
and $\mathbf{e}_{i}$ is the standard basis vector in $\mathbb{R}^{n}$.
Observe that $\mathbf{X}^{i,j}$ are independent, have mean zero, and are also independent of $\mathbf{Y}$, which itself has mean zero. Then, applying the matrix Bernstein inequality (Theorem 5.4.1 in Vershynin (2018)), we obtain:
$\operatorname{\mathbb{P}}\left\{\left\lVert\mathbf{Y}+\sum_{1\leq i<j\leq n}\mathbf{X}^{i,j}\right\rVert\geq t\right\}\leq 2n\exp\left(-\frac{t^{2}/2}{\sigma^{2}+Kt/3}\right),$ (46)
where
$\|\mathbf{Y}\|,\left\lVert\mathbf{X}^{i,j}\right\rVert\leq K=1,$
and
$\sigma^{2}=\left\lVert{\rm I}\kern-1.79993pt{\rm E}\mathbf{Y}^{2}+\sum_{1\leq i<j\leq n}{\rm I}\kern-1.79993pt{\rm E}(\mathbf{X}^{i,j})^{2}\right\rVert.$
Since
$(\mathbf{X}^{i,j})^{2}=(A_{ij}-R_{ij})^{2}(\mathbf{e}_{i}\mathbf{e}_{i}^{\top}+\mathbf{e}_{j}\mathbf{e}_{j}^{\top}),$
we obtain
${\rm I}\kern-1.79993pt{\rm E}(\mathbf{X}^{i,j})^{2}=(R_{ij}-R_{ij}^{2})(\mathbf{e}_{i}\mathbf{e}_{i}^{\top}+\mathbf{e}_{j}\mathbf{e}_{j}^{\top}).$
Similarly,
${\rm I}\kern-1.79993pt{\rm E}\mathbf{Y}^{2}=\sum_{i=1}^{n}(R_{ii}-R_{ii}^{2})\mathbf{e}_{i}\mathbf{e}_{i}^{\top}.$
Since each $A_{ij}$ follows a Bernoulli distribution, we have $R_{ij}\in[0,1]$, and thus, $R_{ij}-R_{ij}^{2}\leq 1/4$. It follows that
$\sigma^{2}\leq n/4.$
Setting $t=3\sqrt{n\log n}$ in (46), we obtain:
$\operatorname{\mathbb{P}}\left\{\left\lVert\sum_{1\leq i<j\leq n}\mathbf{Y}+\mathbf{X}^{i,j}\right\rVert\geq 3\sqrt{n\log n}\right\}$ $\leq 2n\exp\left(-\frac{\frac{9n\log n}{2}}{\frac{n}{4}+\frac{3\sqrt{n\log n}}{3}}\right)$
$\leq 2n\exp(-3\log n)$
$=\frac{2}{n^{2}}.$
$\Box$
Proof of Proposition 5
The proof of Proposition 5 follows from Lemmas 4 to 6. Lemmas 4 and 5 have already been stated; we now proceed to prove them.
Proof of Lemma 4:
We first prove (27). The inequality
$\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))\leq\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))$
follows from the monotonicity property of $\phi$ (Property 1) for $\alpha\geq 0$. For the reverse inequality, we consider
$\log\frac{\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))}{\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))}$ $=\log\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))-\log\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))$
$=(K-1)\min_{i}d_{i}(\mathbf{R})\frac{\partial}{\partial d}\log\phi(\alpha,d_{0}),$
for some $d_{0}\in[\min_{i}d_{i}(\mathbf{R}),K\min_{i}d_{i}(\mathbf{R})]$.
By Property 3 in (13),
$(K-1)\min_{i}d_{i}(\mathbf{R})\frac{\partial}{\partial d}\log\phi(\alpha,d_{0})$ $=(K-1)\min_{i}d_{i}(\mathbf{R})\frac{\partial\phi/\partial d(\alpha,d_{0})}{\phi(\alpha,d_{0})}$
$\leq(K-1)Kd_{0}\frac{\partial\phi/\partial d(\alpha,d_{0})}{\phi(\alpha,d_{0})}\leq C.$
Thus, we obtain
$\frac{\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))}{\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))}=\exp\left(\log\frac{\phi(\alpha,K\min_{i}d_{i}(\mathbf{R}))}{\phi(\alpha,\min_{i}d_{i}(\mathbf{R}))}\right)\leq e^{C}.$
The proof of (28) follows similarly using Property 3 in (14). Finally, (29) and (30) follow from (27) and (28), respectively, by reversing all inequalities, since $\partial\phi/\partial d\leq 0$ for $\alpha<0$.
$\blacksquare$
Proof of Lemma 5:
Recall that
$d_{j}({\mathbf{A}})=\sum_{i=1}^{n}A_{ij},\quad d_{j}(\mathbf{R})=\sum_{i=1}^{n}R_{ij}.$
Applying Hoeffding’s inequality, we obtain
$\operatorname{\mathbb{P}}\{|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})|\geq t\}$ =$=\operatorname{\mathbb{P}}\left\{\left|\sum_{i=1}^{n}(A_{ij}-R_{ij})\right|\geq t\right\}$
$\leq\operatorname{\mathbb{P}}\left\{\sum_{i=1}^{n}(A_{ij}-R_{ij})\geq t\right\}+\operatorname{\mathbb{P}}\left\{\sum_{i=1}^{n}(R_{ij}-A_{ij})\geq t\right\}$
$\leq 2\exp\left(-\frac{2t^{2}}{n}\right).$
Setting $t=\sqrt{2n\log n}$, we obtain
$\operatorname{\mathbb{P}}\{|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})|\geq\sqrt{2n\log n}\}\leq 2e^{-4\log n}=\frac{2}{n^{4}},$ (47)
which establishes (31).
For (32), by the mean value theorem, we have
$\left|\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j}(\mathbf{R}))\right|=\left|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})\right|\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j})\right|,$
for some $c_{j}\in[d_{j}(\mathbf{R})-|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})|,d_{j}(\mathbf{R})+|d_{j}(\mathbf{A})-d_{j}(\mathbf{R})|]$.
Using the bound from (47), it follows that
$\left|\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j}(\mathbf{R}))\right|\leq\sqrt{2n\log n}\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,1})\right|,$
for some $c_{j,1}\in\left[d_{j}(\mathbf{R})-\sqrt{2n\log n},d_{j}(\mathbf{R})+\sqrt{2n\log n}\right]$ for each $j=1,\ldots,n$, with probability at least $1-\frac{2}{n^{4}}$.
$\blacksquare$
Lemma 6.
Under Assumptions (1)-(2),
$\|\mathbf{D}_{\mathbf{A}}-\mathbf{D}_{\mathbf{R}}\|\leq C\sqrt{n\log n},$ (48)
with probability at least $1-\frac{14}{n^{2}}$, where $C$ is a fixed positive constant independent of $n$.
Proof of Lemma 6:
Since both $\mathbf{D}_{\mathbf{A}}$ and $\mathbf{D}_{\mathbf{R}}$ are diagonal matrices, we first derive an entrywise bound. For any $i=1,\ldots,n$, we have
$\left|\frac{\sum_{j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))}{\phi(\alpha,d_{i}(\mathbf{A}))}-\frac{\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))}{\phi(\alpha,d_{i}(\mathbf{R}))}\right|$
$\leq$ $\left|\frac{\sum_{j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))}{\phi(\alpha,d_{i}(\mathbf{A}))}-\frac{\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))}{\phi(\alpha,d_{i}(\mathbf{A}))}\right|+\left|\frac{\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))}{\phi(\alpha,d_{i}(\mathbf{A}))}-\frac{\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))}{\phi(\alpha,d_{i}(\mathbf{R}))}\right|$
[…]
Define $d_{j,-i}(\mathbf{A})=d_{j}(\mathbf{A})-A_{ij}$, noting that $d_{j,-i}(\mathbf{A})$ is independent of $A_{ij}$. Then, the numerator of the first term 1 in (49) can be decomposed as:
$\left|\sum_{j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))-\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))\right|$
$\leq$ $\left|\sum_{j}A_{ij}\left(\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j,-i}(\mathbf{A}))\right)\right|+\left|\sum_{j}(A_{ij}-R_{ij})\phi(\alpha,d_{j,-i}(\mathbf{A}))\right|$
$+\left|\sum_{j}R_{ij}\left[\phi(\alpha,d_{j,-i}(\mathbf{A}))-\phi(\alpha,d_{j,-i}(\mathbf{R}))\right]\right|+\left|\sum_{j}R_{ij}\left[\phi(\alpha,d_{j}(\mathbf{R}))-\phi(\alpha,d_{j,-i}(\mathbf{R}))\right]\right|$
=\,= $\mathbf{G}+\mathbf{H}+\mathbf{J}+\mathbf{K}.$ (50)
Bound for $\mathbf{G}$ in (50):
By the mean value theorem and Lemma 5, we have
|$\left|\left\{\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j,-i}(\mathbf{A}))\right\}-\left\{\phi(\alpha,d_{j}(\mathbf{R}))-\phi(\alpha,d_{j,-i}(\mathbf{R}))\right\}\right|$
$\leq$ |$\,\left|\left\{\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j}(\mathbf{R}))\right\}-\left\{\phi(\alpha,d_{j,-i}(\mathbf{A}))-\phi(\alpha,d_{j,-i}(\mathbf{R}))\right\}\right|$
$\leq$  $\,2\max\left(\left|\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j}(\mathbf{R}))\right|,\left|\phi(\alpha,d_{j,-i}(\mathbf{A}))-\phi(\alpha,d_{j,-i}(\mathbf{R}))\right|\right)$
$\leq$  $\,2\sqrt{2n\log n}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,2})\right|\,,$
for some $c_{j,2}\in\left[d_{j}(\mathbf{R})-\sqrt{2n\log n}-1,d_{j}(\mathbf{R})+\sqrt{2n\log n}\right]$ for each $j=1,\ldots,n$, with probability at least $1-\frac{2}{n^{4}}$.
Applying the triangle inequality,
$\left|\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j,-i}(\mathbf{A}))\right|$ $\leq\left|\phi(\alpha,d_{j}(\mathbf{R}))-\phi(\alpha,d_{j,-i}(\mathbf{R}))\right|+2\sqrt{2n\log n}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,2})\right|$
$\leq\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,3})\right|+2\sqrt{2n\log n}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,2})\right|\,,$
for some $c_{j,3}\in[d_{j,-i}(\mathbf{R}),d_{j}(\mathbf{R})]$ for each $j=1,\ldots,n$, with probability at least $1-\frac{2}{n^{4}}$.
Note that this bound is not optimal but is sufficient for the proof of Lemma 6. Consider the event:
$\Lambda=\bigcap_{j}\left\{\left|\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j,-i}(\mathbf{A}))\right|\leq\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,3})\right|+2\sqrt{2n\log n}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,2})\right|\right\}.$
By the union bound [^29] For events $E_{1},\ldots,E_{n}$ satisfying $\operatorname{\mathbb{P}}(E_{j})\geq 1-c_{j}$ for all $j=1,\ldots,n$, it holds that $\operatorname{\mathbb{P}}(\cap_{j}E_{j})=\operatorname{\mathbb{P}}((\cup_{j}E_{j}^{C})^{C})=1-\operatorname{\mathbb{P}}(\cup_{j}E_{j}^{C})\geq 1-\sum_{j}\operatorname{\mathbb{P}}(E_{j}^{C})\geq 1-\sum_{j}c_{j}$., we obtain $\operatorname{\mathbb{P}}\{\Lambda\}\geq 1-\frac{2}{n^{3}}$.
Denote $\operatorname{\mathbb{P}}_{S}\{T\}=\operatorname{\mathbb{P}}\{S\cap T\}$ for events $S$ and $T$. Define
$t=2\left(\log n\sum_{j}\left(\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,3})\right|+2\sqrt{2n\log n}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,2})\right|\right)^{2}\right)^{1/2}.$ (51)
Applying Hoeffding’s inequality, we obtain
$\operatorname{\mathbb{P}}\left\{\left|\sum_{j}(A_{ij}-R_{ij})\left(\phi(\alpha,d_{j}(\mathbf{A}))-\phi(\alpha,d_{j,-i}(\mathbf{A}))\right)\right|\geq t\right\}$
[…]
Bound for $\mathbf{H}$ in (50):
By the mean value theorem and Lemma 5, we obtain
$\left|\phi(\alpha,d_{j,-i}(\mathbf{A}))-\phi(\alpha,d_{j,-i}(\mathbf{R}))\right|$ $\leq\sqrt{2(n-1)\log(n-1)}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,4})\right|$
$\leq\sqrt{2n\log n}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,4})\right|\,,$
for some $c_{j,4}\in\left[d_{j}(\mathbf{R})-\sqrt{2n\log n}-1,d_{j}(\mathbf{R})+\sqrt{2n\log n}\right]$ for each $j=1,\ldots,n$, with probability at least $1-\frac{2}{n^{4}}$.
Consider the event:
$\Gamma=\bigcap_{j}\left\{\phi(\alpha,d_{j,-i}(\mathbf{A}))\leq\phi(\alpha,d_{j}(\mathbf{R}))+\sqrt{2n\log n}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,4})\right|\right\}.$ (53)
Using the same proof technique as in Lemma 5 and applying the union bound, we obtain $\operatorname{\mathbb{P}}\{\Gamma\}\geq 1-\frac{2}{n^{3}}$.
Define
$t=2\left(\log n\sum_{j}\left(\phi(\alpha,d_{j}(\mathbf{R}))+\sqrt{2n\log n}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,4})\right|\right)^{2}\right)^{1/2}.$ (54)
Applying Hoeffding’s inequality, we obtain
$\operatorname{\mathbb{P}}\left\{\left|\sum_{j}(A_{ij}-R_{ij})\phi(\alpha,d_{j,-i}(\mathbf{A}))\right|\geq t\right\}$
[…]
=\,= $\frac{4}{n^{3}}.$ (55)
Bound for $\mathbf{J}$ in (50):
Conditioning on the event $\Gamma$, we obtain
$\left|\sum_{j}R_{ij}\left[\phi(\alpha,d_{j,-i}(\mathbf{A}))-\phi(\alpha,d_{j,-i}(\mathbf{R}))\right]\right|\leq\sqrt{2n\log n}\sum_{j}R_{ij}\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,4})\right|.$ (56)
Thus, this inequality holds with probability at least $1-\frac{2}{n^{3}}$.
Bound for $\mathbf{K}$ in (50):
By the mean value theorem, we have
$\left|\sum_{j}R_{ij}\left[\phi(\alpha,d_{j}(\mathbf{R}))-\phi(\alpha,d_{j,-i}(\mathbf{R}))\right]\right|\leq\sum_{j}R_{ij}R_{ij}\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,5})\right|\leq\sum_{j}R_{ij}\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,5})\right|,$ (57)
for some $c_{j,5}\in[d_{j,-i}(\mathbf{R}),d_{j}(\mathbf{R})]$ for each $j=1,\ldots,n$.
Bound for 1 in (49):
Assumptions 1 and 2 imply that $d_{j}(\mathbf{R})$ is of the same order for all $j$. By Lemma 4, this also holds for $\phi(\alpha,d_{j}(\mathbf{R}))$ and $\frac{\partial\phi}{\partial d}(\alpha,d_{j}(\mathbf{R}))$. Using assumption 1 and Property 3 (13) of $\phi$, we combine the four bounds (52) – (57) to obtain a bound for the numerator of 1:
$\left|\sum_{j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))-\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))\right|\leq C\sqrt{n\log n}\,\phi(\alpha,d_{i}(\mathbf{R})),$ (58)
with probability at least $1-\frac{10}{n^{3}}$, where $C$ is a fixed constant independent of $n$.
By Lemma 5 and the triangle inequality, we obtain
$\frac{1}{\phi(\alpha,d_{i}(\mathbf{A}))}$ $\leq\frac{1}{\phi(\alpha,d_{i}(\mathbf{R}))-|\phi(\alpha,d_{i}(\mathbf{A}))-\phi(\alpha,d_{i}(\mathbf{R}))|}$
$\leq\frac{1}{\phi(\alpha,d_{i}(\mathbf{R}))-\sqrt{2n\log n}\,\left|\frac{\partial\phi}{\partial d}(\alpha,c_{j,1})\right|},$
with probability at least $1-\frac{2}{n^{4}}$. Therefore, we obtain
[…]
with probability at least $1-\frac{12}{n^{3}}$.
Bound for 2 in (49):
By Lemma 5 and the bound for $\frac{1}{\phi(\alpha,d_{i}(\mathbf{A}))}$, we have
[…]
=\,= $\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))\frac{\left|\phi(\alpha,d_{i}(\mathbf{A}))-\phi(\alpha,d_{i}(\mathbf{R}))\right|}{\phi(\alpha,d_{i}(\mathbf{A}))\phi(\alpha,d_{i}(\mathbf{R}))}$
[…]
with probability at least $1-\frac{2}{n^{4}}$.
[…]
Combining (58) and (59), we obtain the entrywise bound. For any $i$,
$\left|\frac{\sum_{j}A_{ij}\,\phi(\alpha,d_{j}(\mathbf{A}))}{\phi(\alpha,d_{i}(\mathbf{A}))}-\frac{\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))}{\phi(\alpha,d_{j}(\mathbf{R}))}\right|\leq C\sqrt{n\log n},$ (60)
with probability at least $1-\frac{14}{n^{3}}$. Finally, applying the union bound,
$\|\mathbf{D}_{\mathbf{A}}-\mathbf{D}_{\mathbf{R}}\|\leq C\sqrt{n\log n},$ (61)
with probability at least $1-\frac{14}{n^{2}}$.
$\blacksquare$
Proof of Proposition 5
We are now ready to prove Proposition 5. Rewrite the left hand side as:
$\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|$ $=\left\lVert\left(\mathbf{D}_{\mathbf{A}}^{-1}-\mathbf{D}_{\mathbf{R}}^{-1}\right)\left(\mathbf{D}_{\mathbf{A}}^{-1/2}+\mathbf{D}_{\mathbf{R}}^{-1/2}\right)^{-1}\right\rVert$
$=\left\lVert\left(\mathbf{D}_{\mathbf{A}}-\mathbf{D}_{\mathbf{R}}\right)\left(\mathbf{D}_{\mathbf{A}}\mathbf{D}_{\mathbf{R}}\right)^{-1}\left(\mathbf{D}_{\mathbf{A}}^{-1/2}+\mathbf{D}_{\mathbf{R}}^{-1/2}\right)^{-1}\right\rVert$
$\leq\|\mathbf{D}_{\mathbf{A}}-\mathbf{D}_{\mathbf{R}}\|\cdot\left\|\mathbf{D}_{\mathbf{A}}^{-1}\right\|\cdot\left\|\mathbf{D}_{\mathbf{R}}^{-1}\right\|\cdot\left\|\left(\mathbf{D}_{\mathbf{A}}^{-1/2}+\mathbf{D}_{\mathbf{R}}^{-1/2}\right)^{-1}\right\rVert\,.$
Condition on the event $\|\mathbf{D}_{\mathbf{A}}-\mathbf{D}_{\mathbf{R}}\|\leq C\sqrt{n\log n}\,,$
$\left\|\left(\mathbf{D}_{\mathbf{A}}^{-1/2}+\mathbf{D}_{\mathbf{R}}^{-1/2}\right)^{-1}\right\rVert$ $=\max_{i}\left((\mathbf{D}_{\mathbf{A}})^{-1/2}_{ii}+(\mathbf{D}_{\mathbf{R}})^{-1/2}_{ii}\right)^{-1}$
$=\left(\min_{i}(\mathbf{D}_{\mathbf{A}})^{-1/2}_{ii}+(\mathbf{D}_{\mathbf{R}})^{-1/2}_{ii}\right)^{-1}$
$\leq\left(\min_{i}(\mathbf{D}_{\mathbf{R}})^{-1/2}_{ii}+((\mathbf{D}_{\mathbf{R}})_{ii}+|(\mathbf{D}_{\mathbf{A}})_{ii}-(\mathbf{D}_{\mathbf{R}})_{ii}|)^{-1/2}\right)^{-1}$
$\leq C\left(\min_{i}(\mathbf{D}_{\mathbf{R}})^{-1/2}_{ii}\right)^{-1}$
$\leq C\max_{i}\sqrt{d_{i}(\mathbf{R})}$
$\leq C\min_{i}\sqrt{d_{i}(\mathbf{R})}\,.$
Similarly,
$\left\|\mathbf{D}_{\mathbf{A}}^{-1}\right\|=\left(\min_{i}(\mathbf{D}_{\mathbf{A}})_{ii}\right)^{-1}\leq C\left(\min_{i}d_{i}(\mathbf{R})\right)^{-1}\,.$
Combining the results and applying Lemma 6, we have
$\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|\leq C\sqrt{n\log n}\left(\min_{i}d_{i}(\mathbf{R})\right)^{-3/2}\,,$
with probability at least $1-\frac{14}{n^{2}}$. We are now ready to prove Proposition 5. We begin by rewriting the left-hand side as:
$\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|$ $=\left\lVert\left(\mathbf{D}_{\mathbf{A}}^{-1}-\mathbf{D}_{\mathbf{R}}^{-1}\right)\left(\mathbf{D}_{\mathbf{A}}^{-1/2}+\mathbf{D}_{\mathbf{R}}^{-1/2}\right)^{-1}\right\rVert$
$=\left\lVert\left(\mathbf{D}_{\mathbf{A}}-\mathbf{D}_{\mathbf{R}}\right)\left(\mathbf{D}_{\mathbf{A}}\mathbf{D}_{\mathbf{R}}\right)^{-1}\left(\mathbf{D}_{\mathbf{A}}^{-1/2}+\mathbf{D}_{\mathbf{R}}^{-1/2}\right)^{-1}\right\rVert$
$\leq\|\mathbf{D}_{\mathbf{A}}-\mathbf{D}_{\mathbf{R}}\|\cdot\left\|\mathbf{D}_{\mathbf{A}}^{-1}\right\|\cdot\left\|\mathbf{D}_{\mathbf{R}}^{-1}\right\|\cdot\left\|\left(\mathbf{D}_{\mathbf{A}}^{-1/2}+\mathbf{D}_{\mathbf{R}}^{-1/2}\right)^{-1}\right\rVert.$
Conditioning on the event $\|\mathbf{D}_{\mathbf{A}}-\mathbf{D}_{\mathbf{R}}\|\leq C\sqrt{n\log n}$, we bound the inverse term:
$\left\|\left(\mathbf{D}_{\mathbf{A}}^{-1/2}+\mathbf{D}_{\mathbf{R}}^{-1/2}\right)^{-1}\right\rVert$ $=\max_{i}\left((\mathbf{D}_{\mathbf{A}})^{-1/2}_{ii}+(\mathbf{D}_{\mathbf{R}})^{-1/2}_{ii}\right)^{-1}$
$=\left(\min_{i}(\mathbf{D}_{\mathbf{A}})^{-1/2}_{ii}+(\mathbf{D}_{\mathbf{R}})^{-1/2}_{ii}\right)^{-1}$
$\leq\left(\min_{i}(\mathbf{D}_{\mathbf{R}})^{-1/2}_{ii}+((\mathbf{D}_{\mathbf{R}})_{ii}+|(\mathbf{D}_{\mathbf{A}})_{ii}-(\mathbf{D}_{\mathbf{R}})_{ii}|)^{-1/2}\right)^{-1}$
$\leq C\left(\min_{i}(\mathbf{D}_{\mathbf{R}})^{-1/2}_{ii}\right)^{-1}$
$\leq C\max_{i}\sqrt{d_{i}(\mathbf{R})}$
$\leq C\min_{i}\sqrt{d_{i}(\mathbf{R})}.$
Similarly, we bound $\|\mathbf{D}_{\mathbf{A}}^{-1}\|$ as follows:
$\left\|\mathbf{D}_{\mathbf{A}}^{-1}\right\|=\left(\min_{i}(\mathbf{D}_{\mathbf{A}})_{ii}\right)^{-1}\leq C\left(\min_{i}d_{i}(\mathbf{R})\right)^{-1}.$ (62)
Combining these results and applying Lemma 6, we obtain:
$\left\|\mathbf{D}_{\mathbf{A}}^{-1/2}-\mathbf{D}_{\mathbf{R}}^{-1/2}\right\|\leq C\sqrt{n\log n}\left(\min_{i}d_{i}(\mathbf{R})\right)^{-3/2},$ (63)
with probability at least $1-\frac{14}{n^{2}}$. $\Box$
Proof of Remark 1
The proof is essentially the same as that of Proposition 1 and Theorem 1. The main difference lies in controlling $|d_{i}^{\alpha_{i}}-d_{i}^{\alpha_{i}}(\mathbf{R})|$ for varying $\alpha_{i}$. Since $\max_{i,j}|\alpha_{i}-\alpha_{j}|\leq c$, the values of $\alpha_{i}$ concentrate around $\bar{\alpha}=\frac{1}{n}\sum_{i=1}^{n}\alpha_{i}$, implying that $|d_{i}^{\alpha_{i}}-d_{i}^{\alpha_{i}}(\mathbf{R})|$ is of the same order as $|d_{i}^{\bar{\alpha}}-d_{i}^{\bar{\alpha}}(\mathbf{R})|$.
$\Box$
Appendix C Convergence Speed: General Setting and Related Proofs
High-Degree and Low-Degree Model
As in a general stochastic block model, suppose there are $n$ agents in total, divided into $m\geq 2$ groups. Each group has size $n_{i}$ for $i\in\{1,\dots,m\}$. The linking probability matrix $\mathbf{P}$ is given by:
$\mathbf{P}=\begin{bmatrix}p&q&\ldots&q\\ q&p&\ldots&q\\ \vdots&\vdots&\ddots&\vdots\\ q&q&\ldots&p\end{bmatrix}\,,$ (64)
where $p$ is the within-group linking probability, and $q$ is the probability of linking across different groups. The relative size of $p$ and $q$ serves as a measure of homophily; a larger $p$ relative to $q$ indicates a more homophilous network.
To fully characterize the second-largest eigenvalue, we assume that there are only two distinct group sizes:
Assumption 3 (Two Group Sizes).
$n_{1}\neq n_{2}=\ldots=n_{m}.$ (65)
Under these specifications, there are only two types of agents that differ in terms of expected degree (i.e., the expected number of connections):
Lemma 7 (High-Degree and Low-Degree Groups).
Let $d_{i}^{*}$ denote the expected degree of agents in group $i\in\{1,\dots,m\}$. Then, the expected degree of an agent in group $1$ is $d_{1}^{*}=n_{1}p+(m-1)n_{2}q$, and for an agent in any other group, it is $d_{2}^{*}=n_{1}q+n_{2}p+(m-2)n_{2}q$.
We refer to the group(s) with the larger expected degree as high-degree groups and those with the smaller expected degree as low-degree groups.
This simple structure captures the following three cases:
  1. Case 1
    One low-degree group and one high-degree group.
  2. Case 2
    One low-degree group and multiple high-degree groups, reflecting cases where high-degree agents represent different campaigns or hold conflicting views.
  3. Case 3
    One high-degree group and multiple low-degree groups, representing scenarios where multiple (cultural, ethnic, or ideological) groups have divergent initial beliefs.
We now characterize the speed of convergence, which is negatively related to $|\lambda_{2}|$, for each of these three cases.
Proposition 6 ($\lambda_{2}$ in the High-Degree and Low-Degree Model).
Recall that the expected degrees are given by $d_{1}^{*}=n_{1}p+(m-1)n_{2}q$ and $d_{2}^{*}=n_{1}q+n_{2}p+(m-2)n_{2}q$. Let $g:\mathbb{R}\to\mathbb{R}$ be the function:
$g(\alpha)=\frac{\phi(\alpha,d_{2}^{*})}{\phi(\alpha,d_{1}^{*})},$ (66)
where $\phi$ satisfies Properties 1 - 3. Then, under the assumptions on linking probabilities and group sizes in (64) and (65), we have:
  • When $m=2$ (Case 1), the second-largest eigenvalue (in magnitude) is:
    $|\lambda_{2}(\mathbf{T}^{*})|=\left|\frac{n_{1}p\,\phi(\alpha,d_{1}^{*})}{n_{1}p\,\phi(\alpha,d_{1}^{*})+n_{2}q\,\phi(\alpha,d_{2}^{*})}-\frac{n_{1}q\,\phi(\alpha,d_{1}^{*})}{n_{1}q\,\phi(\alpha,d_{1}^{*})+n_{2}p\phi(\alpha,d_{2}^{*})}\right|.$ (67)
  • When $m\geq 3$ and $d_{1}^{*}<d_{2}^{*}$ (Case 2), the second-largest eigenvalue is:
    $|\lambda_{2}(\mathbf{T}^{*})|=\begin{cases}\left|\frac{n_{1}p\,\phi(\alpha,d_{1}^{*})}{n_{1}p\,\phi(\alpha,d_{1}^{*})+(m-1)n_{2}q\,\phi(\alpha,d_{2}^{*})}-\frac{n_{1}q\,\phi(\alpha,d_{1}^{*})}{n_{1}q\,\phi(\alpha,d_{1}^{*})+(n_{2}p+(m-2)n_{2}q)\phi(\alpha,d_{2}^{*})}\right|,&\alpha\leq g^{-1}(n_{1}/n_{2})\\ \left|\frac{n_{2}(p-q)\,\phi(\alpha,d_{2}^{*})}{n_{1}q\,\phi(\alpha,d_{1}^{*})+(n_{2}p+(m-2)n_{2}q)\phi(\alpha,d_{2}^{*})}\right|,&\alpha>g^{-1}(n_{1}/n_{2})\end{cases}$ (68)
  • When $m\geq 3$ and $d_{1}^{*}>d_{2}^{*}$ (Case 3), we have:
    $|\lambda_{2}(\mathbf{T}^{*})|=\begin{cases}\left|\frac{n_{1}p\,\phi(\alpha,d_{1}^{*})}{n_{1}p\,\phi(\alpha,d_{1}^{*})+(m-1)n_{2}q\,\phi(\alpha,d_{2}^{*})}-\frac{n_{1}q\,\phi(\alpha,d_{1}^{*})}{n_{1}q\,\phi(\alpha,d_{1}^{*})+(n_{2}p+(m-2)n_{2}q)\phi(\alpha,d_{2}^{*})}\right|,&\alpha\geq g^{-1}(n_{1}/n_{2})\\ \left|\frac{n_{2}(p-q)\,\phi(\alpha,d_{2}^{*})}{n_{1}q\,\phi(\alpha,d_{1}^{*})+(n_{2}p+(m-2)n_{2}q)\phi(\alpha,d_{2}^{*})}\right|,&\alpha<g^{-1}(n_{1}/n_{2})\end{cases}$ (69)
Proof.
Before proving Proposition 6, we establish the following lemma.
Lemma 8.
Let the $m\times m$ matrix $\mathbf{F}_{\mathbf{T}^{*}}$ be defined as:
$\left(\mathbf{F}_{\mathbf{T}^{*}}\right)_{kl}=\frac{n_{l}P_{kl}\,\phi(\alpha,\sum_{h}n_{h}P_{lh})}{\sum_{l}n_{l}P_{kl}\,\phi(\alpha,\sum_{h}n_{h}P_{lh})}\,.$ (70)
Then, $\mathbf{F}_{\mathbf{T}^{*}}$ has the same eigenvalues as $\mathbf{T}^{*}$.
Proof.
First, note that $\mathbf{R}$ is a matrix with $m^{2}$ blocks. Within each block, the entries are identical, as is $\mathbf{T}^{*}$. From the definition of $\textbf{T}^{*}$, we see that if vertex $i$ belongs to group $k$ and vertex $j$ belongs to group $l$,
$T^{*}_{ij}=\frac{R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))}{\sum_{j}R_{ij}\,\phi(\alpha,d_{j}(\mathbf{R}))}=\frac{P_{kl}\,\phi(\alpha,\sum_{h}n_{h}P_{lh})}{\sum_{l}n_{l}P_{kl}\,\phi(\alpha,\sum_{h}n_{h}P_{lh})}\,.$
After a suitable rearrangement of the vertices, the matrix $\mathbf{T}^{*}$ takes the following block form:
$\begin{bmatrix}\mathbf{B}_{11}&\mathbf{B}_{12}&\ldots&\mathbf{B}_{1m}\\ \mathbf{B}_{21}&\mathbf{B}_{22}&\ldots&\mathbf{B}_{2m}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{B}_{m1}&\mathbf{B}_{m2}&\ldots&\mathbf{B}_{mm}\end{bmatrix}\,,$
where each $\mathbf{B}_{kl}$ is a block matrix in which all entries are identical. Denote the entries in block $\mathbf{B}_{kl}$ by $B_{kl}$, which takes the value of $T_{ij}^{*}$ in (70) if vertex $i$ is in group $k$ and vertex $j$ is in group $l$. Consider eigenvectors in the form of $\mathbf{v}=(v_{1},v_{1},\ldots,v_{2},v_{2},\ldots,v_{m},v_{m}\ldots)^{\top}$. The equation $\mathbf{T}^{*}\mathbf{v}=\lambda\mathbf{v}$ implies:
$n_{1}B_{11}v_{1}+n_{2}B_{12}v_{2}+\ldots+n_{m}B_{1m}v_{m}=\lambda v_{1}\,,$
$\vdots$
$n_{1}B_{m1}v_{1}+n_{2}B_{m2}v_{2}+\ldots+n_{m}B_{mm}v_{m}=\lambda v_{m}\,,$
which completes the proof of Lemma 8. ∎
We focus on the case where $d_{1}^{*}>d_{2}^{*}$ in this proof of Proposition 6. The proof for $d_{1}^{*}<d_{2}^{*}$ is almost identical and thus omitted. By applying Lemma 8, we reduce the $n\times n$ matrix $\mathbf{T}^{*}$ to an $m\times m$ matrix $\mathbf{F_{\mathbf{T}^{*}}}$. Given our assumptions on $\mathbf{n}$ and $\mathbf{P}$ in (65) and (64), there are two expected degrees, denoted as $d_{1}^{*}$ and $d_{2}^{*}$. For vertices in group 1 (the group of size $n_{1}$), $d_{1}^{*}=n_{1}p+(m-1)n_{2}q$, and for vertices in the remaining groups, $d_{2}^{*}=n_{1}q+n_{2}p+(m-2)n_{2}q$. Then, the matrix $\mathbf{F_{\mathbf{T}^{*}}}$ takes the following form:
$\mathbf{F_{\mathbf{T}^{*}}}=\begin{bmatrix}a&b&b&b&\ldots&b\\ e&c&d&d&\ldots&d\\ e&d&c&d&\ldots&d\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ e&d&d&\ldots&c&d\\ e&d&d&\ldots&d&c\\ \end{bmatrix}\,,$ (71)
where
$a=\frac{n_{1}p\,\phi(\alpha,d_{1}^{*})}{n_{1}p\,\phi(\alpha,d_{1}^{*})+(m-1)n_{2}q\,\phi(\alpha,d_{2}^{*})}\,,$
$b=\frac{n_{2}q\,\phi(\alpha,d_{2}^{*})}{n_{1}p\,\phi(\alpha,d_{1}^{*})+(m-1)n_{2}q\,\phi(\alpha,d_{2}^{*})}\,,$
$c=\frac{n_{2}p\,\phi(\alpha,d_{2}^{*})}{n_{1}q\,\phi(\alpha,d_{1}^{*})+n_{2}p\,\phi(\alpha,d_{2}^{*})+(m-2)n_{2}q\,\phi(\alpha,d_{2}^{*})}\,,$
$d=\frac{n_{2}q\,\phi(\alpha,d_{2}^{*})}{n_{1}q\,\phi(\alpha,d_{1}^{*})+n_{2}p\,\phi(\alpha,d_{2}^{*})+(m-2)n_{2}q\,\phi(\alpha,d_{2}^{*})}\,,$
$e=\frac{n_{1}q\,\phi(\alpha,d_{1}^{*})}{n_{1}q\,\phi(\alpha,d_{1}^{*})+n_{2}p\,\phi(\alpha,d_{2}^{*})+(m-2)n_{2}q\,\phi(\alpha,d_{2}^{*})}\,.$
We perform row operations on the matrix $\mathbf{F_{\mathbf{T}^{*}}-\lambda\mathbf{I}_{m}}$ before finding the zeros of the characteristic polynomial:
$\det(\mathbf{F_{\mathbf{T}^{*}}-\lambda\mathbf{I}_{m}})=0\,.$ (72)
Subtracting the last row from rows $12,3,\ldots,m-$ 1, equation (72) becomes:
$\det\begin{bmatrix}a-\lambda&b&b&b&\ldots&b\\ 0&c-d-\lambda&0&0&\ldots&-(c-d-\lambda)\\ 0&0&c-d-\lambda&0&\ldots&-(c-d-\lambda)\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&c-d-\lambda&-(c-d-\lambda)\\ e&d&d&\ldots&d&c-\lambda\\ \end{bmatrix}=0\,.$
We now see that one eigenvalue is $\lambda=c-d$, with algebraic multiplicity $m-2$.
Suppose $\lambda\neq c-d$. Multiplying rows $12,3,\ldots,m-$ 1 by $1/(c-d-\lambda)$, we obtain:
$\det\begin{bmatrix}a-\lambda&b&b&b&\ldots&b\\ 0&1&0&0&\ldots&-1\\ 0&0&1&0&\ldots&-1\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&1&-1\\ e&d&d&\ldots&d&c-\lambda\\ \end{bmatrix}=0\,.$
Then, subtracting multiples of rows $12,3,\ldots,m-$ 1 from the first and last row, we have:
$\det\begin{bmatrix}a-\lambda&0&0&0&\ldots&(m-1)b\\ 0&1&0&0&\ldots&-1\\ 0&0&1&0&\ldots&-1\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\ldots&1&-1\\ e&0&0&\ldots&0&c-\lambda+(m-2)d\\ \end{bmatrix}=0\,.$ (73)
Computing the determinant in (73), we see that the eigenvalues different from $c-d$ should satisfy the equation:
$(a-\lambda)(c-\lambda+(m-2)d)-e(m-1)b=0\,.$
Note that whether $m$ is even or odd does not affect the equation. By solving the quadratic equation, we find that the other two eigenvalues are $\lambda=1$ and $\lambda=a-e$. To determine the second largest eigenvalue, we compare $a-e$ and $c-d$, both of which are positive. Thus, in the computation below, $|\lambda_{2}|=\lambda_{2}$, and the absolute value is omitted. Comparing the two candidates $a-e$ and $c-d$, denoting the denominators of $a$ and $c$ as $D_{1}$ and $D_{2}$ respectively, we obtain
$D_{1}D_{2}(a-e-c+d)$
=\,= $n_{1}p\,\phi(\alpha,d_{1}^{*})[n_{1}q\,\phi(\alpha,d_{1}^{*})+n_{2}p\,\phi(\alpha,d_{2}^{*})+(m-2)n_{2}q\,\phi(\alpha,d_{2}^{*})]$
$-(n_{2}(p-q)\phi(\alpha,d_{2}^{*})+n_{1}q\,\phi(\alpha,d_{1}^{*}))[n_{1}p\,\phi(\alpha,d_{1}^{*})+(m-1)n_{2}q\,\phi(\alpha,d_{2}^{*})]$
=\,= $(m-1)n_{2}q\,\phi(\alpha,d_{2}^{*})(p-q)(n_{1}\,\phi(\alpha,d_{1}^{*})-n_{2}\,\phi(\alpha,d_{2}^{*}))\,.$ (74)
This proves (69) in Proposition 6, since $n_{1}\phi(\alpha,d_{1}^{*})\geq n_{2}\,\phi(\alpha,d_{2}^{*})$ is equivalent to $\alpha\geq g^{-1}(n_{1}/n_{2})$. We are left with the special case of $m=2$. If $m=2$, there is no eigenvalue equal to $c-d$. This completes the computations in Proposition 6.
With $|\lambda_{2}|$ fully characterized, we can now analyze how degree-dependence $\alpha$ affects convergence speed.
Theorem 2 (Non-monotonic Impact of Degree-Dependence, $\alpha$).
Let $\mathcal{D}_{\alpha}$ be the region where $|\lambda_{2}(\mathbf{T}^{*})|$ is decreasing in $\alpha$ for $\alpha\in\mathcal{D}_{\alpha}$ and increasing for $\alpha\notin\mathcal{D}_{\alpha}$. Then, $\mathcal{D}_{\alpha}$ is characterized as follows:
$\mathcal{D}_{\alpha}=\begin{cases}\bigg{[}g^{-1}\left(\frac{n_{1}}{n_{2}}\right),\infty\bigg{)},&\text{if }m=2,\\ (-\infty,g^{-1}(n_{1}/n_{2})]\cup\bigg{[}g^{-1}\left(\frac{n_{1}}{n_{2}}\left(\frac{p}{(m-1)(p+(m-2)q)}\right)^{1/2}\right),\infty\bigg{)},&\text{if }m\geq 3,d_{1}^{*}>d_{2}^{*},\\ \bigg{(}g^{-1}\left(\frac{n_{1}}{n_{2}}\left(\frac{p}{(m-1)(p+(m-2)q)}\right)^{1/2}\right),g^{-1}(n_{1}/n_{2})\bigg{]},&\text{if }m\geq 3,d_{1}^{*}<d_{2}^{*}.\end{cases}$ (75)
Function $g$ is defined by equation (66).
Proof.
In this proof, we focus on the case where $d_{1}^{*}>d_{2}^{*}$. The proof for $d_{1}^{*}<d_{2}^{*}$ follows similarly and is thus omitted. For $m\geq 3$, to show monotonicity, we first consider $\alpha\geq g^{-1}(n_{1}/n_{2})$, for which the second-largest eigenvalue is:
$\lambda_{2}$ $=a-e$
$=\frac{n_{1}p\,\phi(\alpha,d_{1}^{*})}{n_{1}p\,\phi(\alpha,d_{1}^{*})+(m-1)n_{2}q\,\phi(\alpha,d_{2}^{*})}-\frac{n_{1}q\,\phi(\alpha,d_{1}^{*})}{n_{1}q\,\phi(\alpha,d_{1}^{*})+(n_{2}p+(m-2)n_{2}q)\phi(\alpha,d_{2}^{*})}\,,$ (76)
From this expression, we see that the limit is 0 as $\alpha\to\infty$. We rewrite the above expression of $\lambda_{2}$ as:
$\lambda_{2}=\frac{1}{1+C_{1}g(\alpha)}-\frac{1}{1+C_{2}g(\alpha)}\,,$ (77)
where $C_{1}=\frac{(m-1)n_{2}q}{n_{1}p}$, $C_{2}=\frac{n_{2}p+(m-2)n_{2}q}{n_{1}q}$, and $g(\alpha,d_{1}^{*},d_{2}^{*})=\frac{\phi(\alpha,d_{2}^{*})}{\phi(\alpha,d_{1}^{*})}$. Since $d_{1}^{*}$ and $d_{2}^{*}$ are fixed here, we omit them in the notation below and write $g(\alpha)$.
Note that $0<C_{1}<C_{2}$. Then, the derivative with respect to $\alpha$ is:
$\frac{\partial\lambda_{2}}{\partial\alpha}$ $=-\frac{C_{1}\frac{dg}{d\alpha}}{(1+C_{1}g(\alpha))^{2}}+\frac{C_{2}\frac{dg}{d\alpha}}{(1+C_{2}g(\alpha))^{2}}$
$=\frac{dg}{d\alpha}\left[\frac{(C_{2}-C_{1})(1-C_{1}C_{2}g^{2}(\alpha))}{(1+C_{1}g(\alpha))^{2}(1+C_{2}g(\alpha))^{2}}\right]\,.$ (78)
Property 2 in (11) implies that $\frac{dg}{d\alpha}<0$. Thus, $\frac{\partial\lambda_{2}}{\partial\alpha}\leq 0$ for $1-C_{1}C_{2}g^{2}(\alpha)\geq 0$, which is equivalent to:
$\alpha\geq g^{-1}\left(\frac{1}{\sqrt{C_{1}C_{2}}}\right)=g^{-1}\left(\frac{n_{1}}{n_{2}}\left(\frac{p}{(m-1)(p+(m-2)q)}\right)^{1/2}\right)\,.$ (79)
Similarly, for $\alpha<g^{-1}(n_{1}/n_{2})$,
$\lambda_{2}=c-d=\frac{n_{2}(p-q)g(\alpha)}{n_{1}q+n_{2}(p+(m-2)q)g(\alpha)}\,.$ (80)
Then, taking the derivative and simplifying, we get:
$\frac{\partial\lambda_{2}}{\partial\alpha}$ $=\frac{\partial g}{\partial\alpha}\frac{n_{1}n_{2}(p-q)q}{(n_{1}q+n_{2}(p+(m-2)q)g(\alpha))^{2}}<0\,,$ (81)
since $\frac{\partial g}{\partial\alpha}<0$. For $m=2$, noting that $\lambda_{2}=a-e$, the proof follows similarly from (76)–(79).
The remaining part is that although monotonicity in the case of “expectation” does not hold in the case of random networks, we can still determine how much of an increase in $\alpha$ is required for the decrease in the eigenvalue to exceed the error created by randomness. Given an $\alpha_{0}$, suppose $\alpha$ rises from $\alpha_{0}$ to $\alpha_{1}$. How large does $\alpha_{1}$ need to be to observe an increase in convergence speed with high probability? We answer this question by applying Theorem 1 in the following proposition.
Proposition 7.
We assume that $|\alpha_{0}|+|\alpha_{1}|<C$ for some positive constant $C$ and $\alpha_{0},\alpha_{1}\in\mathcal{D}_{\alpha}$ as defined in Theorem 2. There exists an integer $n_{0}\in\mathbb{N}$, such that for all $n>n_{0}$, if
$\alpha_{1}-\alpha_{0}>\hat{C}\frac{\sqrt{\log n}}{\tau_{n}\sqrt{n}}\left[\alpha_{1}d_{1}^{-1}-\alpha_{1}d_{2}^{-1}\right]^{-1},$ (82)
we have
$\operatorname{\mathbb{P}}\left\{|\lambda_{2}(\mathbf{T}(\alpha_{0}))|-|\lambda_{2}(\mathbf{T}(\alpha_{1}))|>0\right\}\geq 1-\frac{16}{n^{2}}\,.$
[…]
Remark 1.
The condition in (82) may seem complicated but is not restrictive: the right-hand side of the inequality approaches zero as $n$ goes to $\infty$. This condition simply describes the minimum required size of $\alpha_{1}-\alpha_{0}$.
This theorem determines the required increase from $\alpha_{0}$ to $\alpha_{1}$ so that the change in the magnitude of the eigenvalue exceeds the error, whose size is given by Theorem 1. Therefore, Proposition 7 states that if $\alpha$ increases from $\alpha_{0}$ to $\alpha_{1}$ satisfying the conditions in Proposition 7, then we will observe a decrease in $|\lambda_{2}({\mathbf{T}})|$ and thus an increase in convergence speed with a probability tending to one as $n$ goes to $\infty$.
Proof.
Recall that in Proposition 6 and Theorem 2, we computed $|\lambda_{2}(\mathbf{T}^{*}(\alpha))|$ and $\frac{\partial|\lambda_{2}(\mathbf{T}^{*}(\alpha))|}{\partial\alpha}$. Since $|\lambda_{2}(\mathbf{T}^{*}(\alpha))|$ is positive and monotonically decreasing on $\mathcal{D}_{\alpha}$, for $\alpha_{1}>\alpha_{0}$,
$|\lambda_{2}(\mathbf{T}^{*}(\alpha_{0}))|-|\lambda_{2}(\mathbf{T}^{*}(\alpha_{1}))|=\lambda_{2}(\mathbf{T}^{*}(\alpha_{0}))-\lambda_{2}(\mathbf{T}^{*}(\alpha_{1}))\,.$ (83)
By the mean value theorem,
$\lambda_{2}(\mathbf{T}^{*}(\alpha_{0}))-\lambda_{2}(\mathbf{T}^{*}(\alpha_{1}))\geq(\alpha_{0}-\alpha_{1})\left.\frac{\partial\lambda_{2}(\mathbf{T}^{*}(\alpha))}{\partial\alpha}\right|_{\alpha=\alpha_{1}}\,.$ (84)
We want to determine the size of the derivative. Let
$C_{1}=\frac{(m-1)n_{2}q}{n_{1}p}\,,$
$C_{2}=\frac{n_{2}p+(m-2)n_{2}q}{n_{1}q}\,.$
Then, for $\alpha\geq g^{-1}(n_{1}/n_{2})$, using the expression (78) in the proof of Theorem 2, we have
$\left.\frac{\partial\lambda_{2}(\mathbf{T}^{*}(\alpha))}{\partial\alpha}\right|_{\alpha=\alpha_{1}}=-C_{3}\left(\frac{n_{2}}{n_{1}},\frac{q}{p},\frac{\phi(\alpha_{1},d_{2})}{\phi(\alpha_{1},d_{1})}\right)\left[\frac{\partial\phi/\partial\alpha(\alpha_{1},d_{1})}{\phi(\alpha_{1},d_{1})}-\frac{\partial\phi/\partial\alpha(\alpha_{1},d_{2})}{\phi(\alpha_{1},d_{2})}\right]\,,$ (85)
where $C_{3}$ is some positive constant that depends on the fractions $\frac{n_{2}}{n_{1}}$, $\frac{q}{p}$, and $\frac{\phi(\alpha_{1},d_{2})}{\phi(\alpha_{1},d_{1})}$, but does not depend on $n$ when $n$ is large. By assumption,
$\alpha_{1}-\alpha_{0}>\hat{C}\left(\frac{n_{2}}{n_{1}},\frac{q}{p},\frac{\phi(\alpha_{1},d_{2})}{\phi(\alpha_{1},d_{1})},\tilde{C}\right)\frac{\sqrt{\log n}}{\tau_{n}\sqrt{n}}\left[\frac{\partial\phi/\partial\alpha(\alpha_{1},d_{1})}{\phi(\alpha_{1},d_{1})}-\frac{\partial\phi/\partial\alpha(\alpha_{1},d_{2})}{\phi(\alpha_{1},d_{2})}\right]^{-1}\,,$
where $\tilde{C}$ is the constant in Theorem 1. Combining (83) – (85), we obtain
$|\lambda_{2}(\mathbf{T}^{*}(\alpha_{0}))|-|\lambda_{2}(\mathbf{T}^{*}(\alpha_{1}))|>\hat{C}C_{3}\frac{\sqrt{\log n}}{\tau_{n}\sqrt{n}}\,,$ (86)
for $\alpha\geq g^{-1}(n_{1}/n_{2})$.
On the other hand, for $\alpha<g^{-1}(n_{1}/n_{2})$, using expression (81) in the proof of Theorem 2, we obtain equation (85) for a positive constant $C_{4}$ possibly different from $C_{3}$. Similar to (86), we then get
$|\lambda_{2}(\mathbf{T}^{*}(\alpha_{0}))|-|\lambda_{2}(\mathbf{T}^{*}(\alpha_{1}))|>\hat{C}C_{4}\frac{\sqrt{\log n}}{\tau_{n}\sqrt{n}}\,,$ (87)
for $\alpha<g^{-1}(n_{1}/n_{2})$.
We then observe that
$|\lambda_{2}(\mathbf{T}(\alpha_{0}))|-|\lambda_{2}(\mathbf{T}(\alpha_{1}))|=\,$ $|\lambda_{2}(\mathbf{T}^{*}(\alpha_{0}))|-|\lambda_{2}(\mathbf{T}^{*}(\alpha_{1}))|+(|\lambda_{2}(\mathbf{T}(\alpha_{0}))|-|\lambda_{2}(\mathbf{T}^{*}(\alpha_{0}))|)$
$\,+(|\lambda_{2}(\mathbf{T}^{*}(\alpha_{1}))|-|\lambda_{2}(\mathbf{T}(\alpha_{1}))|)$
>\,> $|\lambda_{2}(\mathbf{T}^{*}(\alpha_{0}))|-|\lambda_{2}(\mathbf{T}^{*}(\alpha_{1}))|+C_{5}\frac{\sqrt{\log n}}{\tau_{n}\sqrt{n}}\,,$
with probability at least $1-\frac{16}{n^{2}}$ for some constant $C_{5}$, by Theorem 1. Letting $\hat{C}=\frac{(1-C_{5})}{\min(C_{3},C_{4})}$, we see that the result of Theorem 7 follows.
Appendix D Extension: Perturbation of Adjacency Matrix
In this section, we examine the effects of perturbations in the network structure. Specifically, we introduce a perturbation term to each entry of the adjacency matrix $\mathbf{A}$ and analyze its impact on our main results. [^31] Ideally, the perturbation would be applied directly to the matrix $\mathbf{T}$. However, this could easily disrupt the row-stochastic structure of the matrix. For theoretical simplicity, we instead introduce the perturbation to the adjacency matrix $\mathbf{A}$.
More formally, for $i,j=1,2,\ldots,n$, let $\epsilon_{ij}$ be independent and identically distributed random variables supported on $[0,1]$. The perturbed adjacency matrix $\tilde{\mathbf{A}}$ is then defined entrywise as:
$\tilde{A}_{ij}=(1-\delta)A_{ij}+\delta\epsilon_{ij}\,,$
for $\delta\in[0,1]$. Replacing $\mathbf{A}$ with $\tilde{\mathbf{A}}$ in (3), we define the perturbed weight matrix $\tilde{\mathbf{T}}$ by:
$\tilde{T}_{ij}=\frac{\tilde{A}_{ij}\phi(\alpha,d_{j}(\tilde{\mathbf{A}}))}{\sum\limits_{j}\tilde{A}_{ij}\phi(\alpha,d_{j}(\tilde{\mathbf{A}}))}\,.$
Similarly, let $\tilde{\mathbf{R}}={\rm I}\kern-1.79993pt{\rm E}\tilde{\mathbf{A}}$ and define the corresponding deterministic matrix $\tilde{\mathbf{T}}^{*}$ as:
$\tilde{T}^{*}_{ij}=\frac{\tilde{R}_{ij}\phi(\alpha,d_{j}(\tilde{\mathbf{R}}))}{\sum\limits_{j}\tilde{R}_{ij}\phi(\alpha,d_{j}(\tilde{\mathbf{R}}))}\,.$
By similar Proof as Theorem 1 and Bauer-Fike theorem, We then have the following result:
Proposition 8.
Given any degree-dependence parameter $\alpha\in\mathbb{R}$ and a random network $\mathbf{A}(\mathbf{P},\mathbf{n})$, there exists a positive constant $\tilde{C}$, independent of $n$, such that for all $n>0$, the following results hold for the realized learning matrices $\mathbf{T}(\mathbf{A},\alpha)$ and the perturbation learning matrix $\widetilde{\mathbf{T}}=\mathbf{T}(\widetilde{\mathbf{A}},\alpha)$:
$\operatorname{\mathbb{P}}\Bigl{\{}\bigl{|}\lambda_{2}(\mathbf{T})-\lambda_{2}(\widetilde{\mathbf{T}})\bigr{|}\geq\tilde{C}\frac{\sqrt{n\log n}}{\min_{i}d_{i}(\mathbf{R})}\Bigr{\}}\leq\frac{16}{n^{2}}\,.$
The additional perturbation term does not alter the structural assumptions imposed in the paper. Moreover, the concentration result of Theorem 1 continues to hold, as the perturbation $\epsilon_{ij}$ is bounded. Consequently, the concentration inequalities used in the proof of Theorem 1 remain unaffected, ensuring that the results that rely on this theorem follow.